### v1

# Real enumerative invariants relative to the toric boundary and their refinement

Abstract
We introduce new invariants of the projective plane (and, more generally, of certain toric surfaces) that arise from appropriate enumeration of real curves of genus one and two. These invariants admit a refinement (according to the quantum index) similar to the one introduced by Grigory Mikhalkin in the rational case.
MSC-2010 classification: Primary 14N10, Secondary 14J26, 14P05, 14T90
1 Introduction
Refined enumerative geometry, initiated in [1, 7], became one of the central topics in enumerative geometry with important links to closed and open Gromov-Witten invariants and to Donaldson-Thomas invariants. In a big part of known examples, refined invariants appear as one-parameter deformations of complex enumerative invariants (see, for example, [1, 4, 6, 7]). In his groundbreaking paper [19], G. Mikhalkin has proposed a refined invariant provided by enumeration of real rational curves and has related this invariant to the refined tropical invariants of F. Block and L. Göttsche [1]. Namely, he introduced an integer-valued quantum index for real algebraic curves in toric surfaces. To have a quantum index, a real curve should satisfy certain assumptions: it has to intersect toric divisors only at real points and to be irreducible and separating; the latter condition means that, in the complex point set of the normalization of the curve, the complement of the real part is disconnected, i.e., formed by two halves exchanged by the complex conjugation (in fact, the quantum index is associated to a half of a separating real curve, while the other half has the opposite quantum index; for detailed definitions, see Section 3.2). Mikhalkin [19] showed that, for an appropriate kind of constraints, a Welschinger-type enumeration of real rational curves (cf. [25]) in a given divisor class and with a given quantum index produces an invariant and can be directly related to the numerator of a Block-Göttsche refined tropical invariant (represented as a fraction with the standard denominator).
The main goal of the present paper is to introduce refined invariants of Mikhalkin’s type in the case of curves of genus $1$ and $2$. We follow the ideas of [23] and choose constraints so that every counted real curve of genus $g=1$ or $2$ appears to be a maximal one (i.e., it has $g+1$ global real branches), and hence is separating. More precisely, given a toric surface with the tautological real structure and a very ample divisor class, we fix maximally many real points in a generic Menelaus position (see Section 2.2) on the toric boundary of the positive quadrant, where genus $g$ curves from the given linear system must be tangent to toric divisors with prescribed even intersection multiplicities, and we fix $g$ more generic real points inside different non-positive quadrants as extra constraints (through which the curves under consideration should pass). There are finitely many real curves of genus $g$ matching the constraints and all these curves are separating. Their halves have quantum index, and we equip each curve with a certain Welschinger-type sign.
For some toric surfaces (including the projective plane; in the case of genus $2$, in order to avoid cumbersome technicalities, this is the only toric surface considered in the paper), we prove that the signed enumeration of (halves of) real curves of genus $g=1$ or $2$ that match given constraints, belong to a given linear system, and have a prescribed quantum index does not depend on the choice of a (generic) position of the constraints; precise statements can be found in Section 3.3, Theorem 3.7, and Section 3.4, Theorem 3.9. The resulting invariants are said to be refined. In particular, we get new real enumerative invariants (without prescribing values for quantum index) in genus one and two. Tropical counterparts of the above refined invariants, their computation and comparison with the Block-Göttsche refined tropical invariants will be presented in the forthcoming papers [16, 17].
The present paper is organized as follows. Section 2 contains preparatory material. In Section 3, we introduce refined invariants arising in enumeration of real curves of genus zero, one and two (the genus zero version slightly generalizes Mikhalkin’s refined rational invariants). Section 4 (respectively, 5, 6) is devoted to a proof of the invariance statement in genus zero (respectively, one, two).
Acknowledgements. We started this work during our stay at the Mittag-Leffler Institute, Stockholm, in April 2018 and during the visit of the second author to the École Normale Supérieure, Paris, in June 2019, and we completed the work during our research stays at the Mathematisches Forschungsinstitut Oberwolfach in July 2021 and March 2022 (in the framework of Research-in-Pairs program) and in February - March 2023 (in the framework of Oberwolfach Research Fellows program). We are very grateful to these institutions for the support and excellent working conditions. The first author was supported in part by the ANR grants ANR-18-CE40-0009 ENUMGEOM and ANR-22-CE40-0014 SINTROP. The second author has been partly supported by the Israeli Science Foundation grant no. 501/18, and by the Bauer-Neuman Chair in Real and Complex Geometry.
2 Preparation
2.1 Convex lattice polygons and real toric surfaces
Consider the lattice ${\mathbb{Z}}^{2}$ and its ambient plane ${\mathbb{R}}^{2}={\mathbb{Z}}^{2}\otimes{\mathbb{R}}$. Let $P\subset{\mathbb{R}}^{2}$ be a non-degenerate convex lattice polygon. For a lattice segment $\sigma\subset{\mathbb{R}}^{2}$ (respectively, a vector ${\boldsymbol{a}}\in{\mathbb{Z}}^{2}\setminus\{(0,0)\}$), denote by $\|\sigma\|_{\mathbb{Z}}$ (respectively, $\|{\boldsymbol{a}}\|_{\mathbb{Z}}$) its lattice length, i.e., the ratio of the Euclidean length and the minimal length of a non-zero parallel integral vector.
Denote by ${\operatorname{Tor}}(P)$ the complex toric surface associated with $P$. Let ${\operatorname{Tor}}(P)^{\times}\simeq({\mathbb{C}}^{2})^{\times}$ (respectively, ${\operatorname{Tor}}(\partial P)$) be the dense orbit (respectively, the union of all toric divisors) of ${\operatorname{Tor}}(P)$. The toric surface ${\operatorname{Tor}}(P)$ has the tautological real structure, and the real part ${\operatorname{Tor}}_{\mathbb{R}}(P)$ of ${\operatorname{Tor}}(P)$ contains the positive quadrant ${\operatorname{Tor}}_{\mathbb{R}}(P)^{+}\simeq({\mathbb{R}}_{>0})^{2}$. The closure ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P)$ of ${\operatorname{Tor}}_{\mathbb{R}}(P)^{+}$ (with respect to the Euclidean topology) is diffeomorphically taken onto $P$ by the moment map. We pull back the standard orientation and the metric of $P\subset{\mathbb{R}}^{2}$ to ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P)$ and induce an orientation and a metric on the boundary $\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P)$ of ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P)$; in particular, we get a cyclic order on the sides of $P$; we call this order positive. Denote by ${\mathcal{L}}_{P}$ the tautological line bundle over ${\operatorname{Tor}}(P)$; the global sections of ${\mathcal{L}}_{P}$ are spanned by the monomials $z^{\omega}$, $z=(z_{1},z_{2})$, $\omega\in P\cap{\mathbb{Z}}^{2}$.
For each edge $\sigma\subset\partial P$, we consider the toric curve ${\operatorname{Tor}}(\sigma)$, its dense orbit ${\operatorname{Tor}}(\sigma)^{\times}$, the real part ${\operatorname{Tor}}_{\mathbb{R}}(\sigma)$ of ${\operatorname{Tor}}(\sigma)$, and the positive half-axis ${\operatorname{Tor}}_{\mathbb{R}}(\sigma)^{+}\subset{\operatorname{Tor}}(\sigma)^{\times}\cap{\operatorname{Tor}}_{\mathbb{R}}(\sigma)$. The closure ${\operatorname{Tor}}_{\mathbb{R}}^{+}(\sigma)$ of ${\operatorname{Tor}}_{\mathbb{R}}(\sigma)^{+}$ coincides with ${\operatorname{Tor}}(\sigma)\cap\partial{\operatorname{Tor}}^{+}_{\mathbb{R}}(P)$.
2.2 Toric degree and Menelaus condition
A multi-set $\Delta\subset{\mathbb{Z}}^{2}\setminus\{(0,0)\}$ is said to be
  • balanced if the sum of vectors in $\Delta$ is equal to 00,
  • non-degenerate if the vectors of $\Delta$ span ${\mathbb{R}}^{2}$,
  • even if $\Delta\subset(2{\mathbb{Z}})^{2}$.
A balanced non-degenerate multi-set $\Delta\subset{\mathbb{Z}}^{2}\setminus\{(0,0)\}$ is called a (toric) degree.
Let $\Delta\subset{\mathbb{Z}}^{2}\setminus\{(0,0)\}$ be a non-degenerate balanced multi-set. For each vector ${\boldsymbol{a}}\in\Delta$, denote by $\check{\boldsymbol{a}}$ the vector obtained from ${\boldsymbol{a}}$ by the counter-clockwise rotation by $\pi/2$. The vectors $\check{\boldsymbol{a}}$ can be attached to each other so that the next vector starts at the end of the preceding one in order to form a simple broken line bounding a convex lattice polygon whose boundary gets the counter-clockwise orientation. The latter polygon is denoted by $P_{\Delta}$; it is determined by $\Delta$ up to translation. For a convex lattice polygon $\delta$, we set ${\mathcal{A}}(\delta)$ to be the Euclidean area of $\delta$ and ${\mathcal{I}}(\delta)$ to be the number of integer points in the interior of $\delta$. In the case of $P_{\Delta}$, we shortly write ${\mathcal{A}}(\Delta)$ and ${\mathcal{I}}(\Delta)$. The elements of $\Delta$ are denoted by
${\boldsymbol{a}}^{\sigma}_{i},\quad i=1,\ldots,n^{\sigma},$
where $\sigma$ ranges over the set $P^{1}_{\Delta}$ of sides of the polygon $P_{\Delta}$, and $n^{\sigma}$ is the number of vectors in $\Delta$ that are outer normals to $\sigma$.
From now on, assume that $\Delta$ is even; for each $\sigma\in P^{1}_{\Delta}$ and each $i=1,\ldots,n^{\sigma}$, put $2k^{\sigma}_{i}=\|{\boldsymbol{a}}^{\sigma}_{i}\|_{\mathbb{Z}}$.
Definition 2.1
For each $\sigma\in P^{1}_{\Delta}$, consider a sequence ${\boldsymbol{z}}^{\sigma}$ of $n^{\sigma}$ points in ${\operatorname{Tor}}(\sigma)^{\times}$, and denote by ${\boldsymbol{z}}$ the double index sequence (the upper index being $\sigma\in P^{1}_{\Delta}$ and the lower index being $i=1,\ldots,n^{\sigma}$) formed by the sequences ${\boldsymbol{z}}^{\sigma}$. We say that the sequence ${\boldsymbol{z}}$ satisfies the Menelaus $\Delta$-condition (cf. [19, Section 5.1]) if there exists a curve $C\in|{\mathcal{L}}_{P_{\Delta}}|$ such that $C$ does not contain toric divisors as components and, for each $\sigma\in P^{1}_{\Delta}$, the scheme-theoretic intersection of $C$ and ${\operatorname{Tor}}(\sigma)$ coincides with $\sum_{i=1}^{n^{\sigma}}2k^{\sigma}_{i}z^{\sigma}_{i}$.
Lemma 2.2
The sequences ${\boldsymbol{z}}$ satisfying the Menelaus $\Delta$-condition form an algebraic hypersurface $\operatorname{Men}(\Delta)\subset\prod_{\sigma\in P^{1}_{\Delta}}{\operatorname{Tor}}(\sigma)^{n^{\sigma}}$.
Proof. We present an explicit equation of $\operatorname{Men}(\Delta)$. Consider a linear functional $\lambda:{\mathbb{R}}^{2}\to{\mathbb{R}}$ which is injective on ${\mathbb{Z}}^{2}$. The maximal and the minimal points of $\lambda$ on $P_{\Delta}$ divide the boundary $\partial P_{\Delta}$ of $P_{\Delta}$ into two broken lines $B^{\prime},B^{\prime\prime}$, and the maximal and minimal points of $\lambda$ on each side $\sigma$ define its orientation. An automorphism of ${\mathbb{Z}}^{2}$, which takes an oriented side $\sigma$ of $P$ onto the naturally oriented segment $[0,\|\sigma\|_{\mathbb{Z}}]$ of the first axis of ${\mathbb{R}}^{2}$, defines an isomorphism of ${\operatorname{Tor}}(\sigma)^{\times}$ with ${\mathbb{C}}^{\times}$. Denote by $\xi^{\sigma}=(\xi_{i}^{\sigma})_{i=1}^{n^{\sigma}}$ the sequence of images in ${\mathbb{C}}^{\times}$ of the points of ${\boldsymbol{z}}^{\sigma}$. Then, $\operatorname{Men}(\Delta)$ is given by the equation
$\prod_{\sigma\subset B^{\prime}}\prod_{i=1}^{n^{\sigma}}(\xi_{i}^{\sigma})^{2k_{i}^{\sigma}}=\prod_{\sigma\subset B^{\prime\prime}}\prod_{i=1}^{n^{\sigma}}(\xi_{i}^{\sigma})^{2k_{i}^{\sigma}}\ .$
$\Box$
The hypersurface $\operatorname{Men}(\Delta)$ is called Menelaus hypersurface. Notice that it is reducible. More precisely, $\operatorname{Men}(\Delta)$ splits into $2k_{0}$ components, where
$k_{0}=\gcd\{k_{i}^{\sigma}\ :\ i=1,...,n^{\sigma},\ \sigma\subset\partial P_{\Delta}\}.$
Denote by $\operatorname{Men}(\Delta)_{red}$ the component given by the equation
$\prod_{\sigma\subset B^{\prime}}\prod_{i=1}^{n^{\sigma}}(\xi_{i}^{\sigma})^{k_{i}^{\sigma}/k_{0}}=\prod_{\sigma\subset B^{\prime\prime}}\prod_{i=1}^{n^{\sigma}}(\xi_{i}^{\sigma})^{k_{i}^{\sigma}/k_{0}}\ .$ (1)
Note that the introduced above isomorphism ${\operatorname{Tor}}(\sigma)^{\times}\simeq{\mathbb{C}}^{\times}$ takes ${\operatorname{Tor}}_{\mathbb{R}}(\sigma)^{+}$ onto ${\mathbb{R}}_{>0}\subset{\mathbb{C}}^{\times}$. For any $\tau>0$, denote by $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ the part of $\operatorname{Men}(\Delta)_{red}$ specified by the condition that, for each $\sigma\in P^{1}_{\Delta}$, all points of the subsequence ${\boldsymbol{w}}^{\sigma}$ belong to ${\operatorname{Tor}}_{\mathbb{R}}^{+}(\sigma)$ and lie at the distance $\geq\tau$ from the endpoints of ${\operatorname{Tor}}_{\mathbb{R}}^{+}(\sigma)$.
Lemma 2.3
For a sufficiently small $\tau>0$, the (metric) closure $\overline{\operatorname{Men}}_{\mathbb{R}}^{\tau}(\Delta)\subset\prod_{\sigma\subset\partial P}({\operatorname{Tor}}_{\mathbb{R}}^{+}(\sigma))^{n^{\sigma}}$ is diffeomorphic to a convex polytope.
Proof. Since $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ is given by equation (1) with positive variables $\xi_{i}^{\sigma}$ satisfying restrictions of the form $0<{\operatorname{const}}_{1}\leq\xi_{i}^{\sigma}\leq{\operatorname{const}}_{2}<\infty$, the coordinate-wise logarithm takes $\overline{\operatorname{Men}}_{\mathbb{R}}^{\tau}(\Delta)$ onto a hyperplane section of a convex polytope. $\Box$
2.3 Curves on toric surfaces
Given a morphism ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ of a curve $\widehat{C}$ to ${\operatorname{Tor}}(P_{\Delta})$, we denote by
  • ${\Phi}_{*}(\widehat{C})$ the scheme-theoretic push-forward, i.e., the one-dimensional part of the image, whose components are taken with the corresponding multiplicities;
  • ${\Phi}(\widehat{C})$ the reduced model of ${\Phi}_{*}(\widehat{C})$, where all components are taken with multiplicity one;
  • ${\Phi}^{*}(D)$, where $D\subset{\operatorname{Tor}}(P_{\Delta})$ is a divisor intersecting ${\Phi}(\widehat{C})$ in a finitely many points, the divisor on $\widehat{C}$ which is the pull-back of the scheme-theoretic intersection $D\cap{\Phi}_{*}(\widehat{C})$.
Denote by ${\mathcal{M}}_{g,m}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$ the space of isomorphism classes of maps ${\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})$, where $\widehat{C}$ is a smooth connected curve of genus $g$ and ${\boldsymbol{p}}$ is a sequence of $m$ distinct points in $\widehat{C}$, such that ${\Phi}_{*}(\widehat{C})\in|{\mathcal{L}}_{P_{\Delta}}|$. Correspondingly, we denote by ¯$\overline{\mathcal{M}}_{g,m}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$ the Kontsevich compactification of the above moduli space; this compactification is obtained by adding isomorphism classes of stable maps ${\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})$, where $\widehat{C}$ is a connected nodal curve of arithmetic genus $g$ and ${\boldsymbol{p}}$ is a sequence of $m$ distinct points in ${\widehat{C}}\setminus{\operatorname{Sing}}\hskip 1.70709pt({\widehat{C}})$, such that ${\Phi}_{*}(\widehat{C})\in|{\mathcal{L}}_{P_{\Delta}}|$ (here ${\operatorname{Sing}}\hskip 1.70709pt({\widehat{C}})$ stands for the set of singularities of $\widehat{C}$). In what follows, we work with certain subspaces of these moduli spaces specified for genus $g=0$ in Section 3.2, for genus $g=1$ in Section 3.3, and for genus $2$ in Section 3.4). For any subset ${\mathcal{M}}\subset{\mathcal{M}}_{g,m}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$, denote by $\overline{\mathcal{M}}$ the closure of $\mathcal{M}$ in the compactified moduli space ¯$\overline{\mathcal{M}}_{g,m}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$.
We denote by $(C_{1}\cdot C_{2})_{z}$ the intersection multiplicity of curves $C_{1},C_{2}\subset{\operatorname{Tor}}(P_{\Delta})$ at a smooth point $z$ of ${\operatorname{Tor}}(P_{\Delta})$. By $C_{1}C_{2}$ we mean the total intersection multiplicity of the curves $C_{1},C_{2}\subset{\operatorname{Tor}}(P_{\Delta})$ provided that all intersection points of these curves are smooth in ${\operatorname{Tor}}(P_{\Delta})$. Given a local branch $B$ of a curve germ $(C,z)\subset{\operatorname{Tor}}(P_{\Delta})$ at a smooth point $z\in{\operatorname{Tor}}(P_{\Delta})$, we denote by ${\operatorname{ord}}B$ the intersection multiplicity of $B$ with a generic smooth curve through $z$.
Recall that deformations of a morphism ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ of a smooth curve $\widehat{C}$ are encoded by the normal sheaf on $\widehat{C}$ (see, for instance, [8]):
${\mathcal{N}}_{\Phi}:={\Phi}^{*}{\mathcal{T}}{\operatorname{Tor}}(P_{\Delta})/{\mathcal{T}}\widehat{C}$
(where ${\mathcal{T}}$ denotes the tangent bundle). In the case of an immersion, ${\mathcal{N}}_{\Phi}$ is a line bundle of degree
$\deg{\mathcal{N}}_{\Phi}=c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})-2.$ (2)
Let $\Delta=({\boldsymbol{a}}^{\sigma}_{i},\ i=1,...,n^{\sigma},\ \sigma\in P_{\Delta}^{1})$ be a toric degree as introduced in Section 2.2. Fix a non-negative integer $g\leq{\mathcal{I}}(\Delta)$ and a non-negative integer $n(\sigma)\leq n^{\sigma}$ for each edge $\sigma\in P^{1}_{\Delta}$ such that $\sum_{\sigma\in P^{1}_{\Delta}}n(\sigma)<\sum_{\sigma\in P^{1}_{\Delta}}n^{\sigma}$. Put $n_{\partial}=\sum_{\sigma\in P^{1}_{\Delta}}n(\sigma)$ and $n_{{\operatorname{in}}}=n-n_{\partial}$, where
$n=\sum_{\sigma\in P^{1}_{\Delta}}n^{\sigma}+g-1.$
Choose a sequence ${\boldsymbol{w}}$ of $n$ distinct points in ${\operatorname{Tor}}(P_{\Delta})$ splitting into two subsequences:
  • ${\boldsymbol{w}}_{\partial}$ consisting of $n_{\partial}$ points in general position on ${\operatorname{Tor}}(\partial P_{\Delta})$ so that, for each side $\sigma\in P^{1}_{\Delta}$, the sequence ${\boldsymbol{w}}_{\partial}$ contains $n(\sigma)$ points ${\boldsymbol{w}}^{\sigma}_{i}$, $1\leq i\leq n(\sigma)$, lying on ${\operatorname{Tor}}(\sigma)$;
  • ${\boldsymbol{w}}_{{\operatorname{in}}}$ consisting of $n_{{\operatorname{in}}}$ points $w_{j}$, $j=1$, $\ldots$, $n_{{\operatorname{in}}}$ in general position in (×$({\mathbb{C}}^{\times})^{2}\subset{\operatorname{Tor}}(P_{\Delta})$.
Introduce the subset ${\mathcal{M}}_{g,n}(\Delta,{\boldsymbol{w}})\subset{\mathcal{M}}_{g,n}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$ consisting of the elements $\left[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})\right]$ satisfying the following assumptions:
  1. (M1)
    the sequence ${\boldsymbol{p}}$ splits into two disjoint subsequences ${\boldsymbol{p}}_{\partial}$ and ${\boldsymbol{p}}_{{\operatorname{in}}}$ containing $n_{\partial}$ points $p^{\sigma}_{i}$, $\sigma\in P^{1}_{\Delta}$, $1\leq i\leq n(\sigma)$, and $n_{{\operatorname{in}}}$ points $p_{j}$, $j=1$, $\ldots$, $n_{{\operatorname{in}}}$, respectively, such that ${\Phi}({\boldsymbol{p}}_{\partial})={\boldsymbol{w}}_{\partial}$ (more precisely, for each $\sigma\in P^{1}_{\Delta}$, one has ${\Phi}(p^{\sigma}_{i})=w^{\sigma}_{i}$, $i=1$, $\ldots$, $n(\sigma)$) and ${\Phi}(p_{j})=w_{j}$, $j=1$, $\ldots$, $n_{{\operatorname{in}}}$;
  2. (M2)
    for each $\sigma\in P^{1}_{\Delta}$, one has
    ${\Phi}^{*}({\operatorname{Tor}}(\sigma))=\sum_{i=1}^{n^{\sigma}}k_{i}^{\sigma}p_{i}^{\sigma}\in{\operatorname{Div}}(\widehat{C}),$
    where ${\Phi}(p_{i}^{\sigma})=w_{i}^{\sigma}$ for all $p^{\sigma}_{i}\in{\boldsymbol{p}}_{\partial}$ and $w^{\sigma}_{i}\in{\boldsymbol{w}}_{\partial}$, $\sigma\in P^{1}_{\Delta}$, $1\leq i\leq n(\sigma)$, and $p^{\sigma}_{i}$, $n(\sigma)<i\leq n^{\sigma}$, are some points of $\widehat{C}$ not counted in the sequence ${\boldsymbol{p}}_{\partial}$.; at last, ${\operatorname{Div}}(\widehat{C})$ denotes the group of Weil divisors on $\widehat{C}$.
Lemma 2.4
Suppose that the sequence ${\boldsymbol{w}}$ is in general position subject to the location with respect to the toric divisors as indicated above. Then, the space ${\mathcal{M}}_{g,n}(\Delta,{\boldsymbol{w}})$ is finite. Moreover, for each element $\left[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})\right]\in{\mathcal{M}}_{g,n}(\Delta,{\boldsymbol{w}})$, the map ${\Phi}$ takes $\widehat{C}$ birationally onto an immersed curve $C={\Phi}(\widehat{C})$ that, for each $\sigma\in P^{1}_{\Delta}$, intersects the toric divisor ${\operatorname{Tor}}(\sigma)$ at $n^{\sigma}$ distinct points and $C$ is smooth at each of these intersection points.
Proof. Consider the family ${\mathcal{M}}^{\prime}_{g,n}(\Delta)$ formed by the elements $[{\Phi}:(\widehat{C},{\boldsymbol{p}}_{\partial})\to{\operatorname{Tor}}(P_{\Delta})]$ having ${\Phi}$ birational onto its image and matching the condition (M2) above. We claim that
$\dim{\mathcal{M}}^{\prime}_{g,n}(\Delta)\leq n.$ (3)
Assuming (3) and imposing the conditions ${\Phi}(p^{\sigma}_{i})=w^{\sigma}_{i}\in{\boldsymbol{w}}_{\partial}$, $1\leq i\leq n(\sigma)$, for every edge $\sigma\in P^{1}_{\Delta}$, as well as imposing extra $n_{{\operatorname{in}}}$ point constraints ${\boldsymbol{w}}_{{\operatorname{in}}}$ in general position in (*$({\mathbb{C}}^{*})^{2}$, we obtain a finite set.
To prove (3), take a generic element $\xi=[{\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})]$ in an irreducible component of ${\mathcal{M}}^{\prime}_{g,n}(\Delta)$. It is a smooth point of that component, and the germ of ${\mathcal{M}}^{\prime}_{g,n}(\Delta)$ at $\xi$ realizes an equisingular deformation. We estimate the dimension of the tangent space to this germ. Recall that, for an isolated singular point $z$ of a planar curve $C$, center of local branches $Q_{1},...,Q_{s}$ (i.e., irreducible components of the germ $(C,z)$), the following holds (see [10, Theorem 1 and formulas (19), (21), (22)]):
  • the tangent space $T(ES_{z})$ to the equisingular stratum $ES_{z}$ in the base space $B_{z}$ of its versal deformation is contained in
    $\left\{b\in B_{z}\ \Big{|}\ \begin{array}[]{c}{\varphi_{b}\in{\mathcal{O}}_{C,z},\ {\operatorname{ord}}_{Q_{i}}\varphi_{b}\geq 2\delta(Q_{i})+\sum_{j\neq i}(Q_{i}\cdot Q_{j})}\\ {\qquad\qquad\qquad\qquad\qquad+{\operatorname{ord}}Q_{i}-1,\quad i=1,...,s}\end{array}\right\};$
  • the tangent space $T(ES_{z}(L))$, where $L\supset\{z\}$ is a line, to the equisingular stratum $ES_{z}(L)$ in the base space $B_{z}(L)$ of a versal deformation constrained by the fixed intersection multiplicity with the line $L$ at the singular point, is contained in
    $\left\{b\in B_{z}\ \Big{|}\ \begin{array}[]{c}{\varphi_{b}\in{\mathcal{O}}_{C,z},\ {\operatorname{ord}}_{Q_{i}}\varphi_{b}\geq 2\delta(Q_{i})+\sum_{j\neq i}(Q_{i}\cdot Q_{j})}\\ {\qquad\qquad\qquad\qquad\qquad\quad+(Q_{i}\cdot L)-1,\quad i=1,...,s}\end{array}\right\};$
  • the tangent space $T(ES^{fix}_{z}(L))$, where $L\supset\{z\}$ is a line, to the equisingular stratum $ES^{fix}_{z}(L)$ in the base space $B^{fix}_{z}(L)$ of a versal deformation constrained by the fixed position of the singular point $z$ and the fixed intersection multiplicity with the line $L$ at $z$, is contained in
    $\left\{b\in B_{z}\ \Big{|}\ \begin{array}[]{c}{\varphi_{b}\in{\mathcal{O}}_{C,z},\ {\operatorname{ord}}_{Q_{i}}\varphi_{b}\geq 2\delta(Q_{i})+\sum_{j\neq i}(Q_{i}\cdot Q_{j})+{\operatorname{ord}}Q_{i}}\\ {\qquad\qquad\qquad\qquad+(Q_{i}\cdot L)-1,\quad i=1,...,s}\end{array}\right\}.$
Let $C={\Phi}(\widehat{C})$. Combining the above description of the ambient space for the considered tangent spaces and taking into account that the conductor ideal $\operatorname{Ann}({\Phi}_{*}{\mathcal{O}}_{\widehat{C}}/{\mathcal{O}}_{C})$ is given by
$\left\{\varphi\in{\mathcal{O}}_{C,z}\ \big{|}\ {\operatorname{ord}}_{Q_{i}}\varphi\geq 2\delta(Q_{i})+\sum_{j\neq i}(Q_{i}\cdot Q_{j}),\ i=1,...,s\right\},$
we derive that
$\dim{\mathcal{M}}^{\prime}_{g,n}(\Delta)\leq h^{0}\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-D)\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{O}}_{{\operatorname{Tor}}(P_{\Delta})}(C)\right),$ (4)
where
$\deg D=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+\sum_{\sigma\in P^{1}_{\Delta}}\sum_{w\in C\cap{\operatorname{Tor}}(\sigma)}\sum_{Q\subset(C,w)}((Q\cdot{\operatorname{Tor}}(\sigma))_{w}-1)$
$+\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap({\mathbb{C}}^{*})^{2}}\sum_{Q\subset(C,z)}({\operatorname{ord}}Q-1).$
Then,
$\deg\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-D)\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{O}}_{{\operatorname{Tor}}(P_{\Delta})}(C)\right)=C^{2}-\deg D=C^{2}-(C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})}+2-2g)$
$+CK_{{\operatorname{Tor}}(P_{\Delta})}+\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}-\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap({\mathbb{C}}^{*})^{2}}\sum_{Q\subset(C,z)}({\operatorname{ord}}Q-1)$
$(2g-2)+\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}-\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap({\mathbb{C}}^{*})^{2}}\sum_{Q\subset(C,z)}({\operatorname{ord}}Q-1).$ (5)
Here,
$3\leq\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}\leq\#\Delta.$ (6)
Then, by (5) and by the Riemann-Roch formula, the right-hand side of (4) does not exceed
$12g-2+\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}-g+$ 1
$=g-1+\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}\leq g-1+\#\Delta=n.$
Furthermore, if either #$\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}<\#\Delta$, or there is a singular local branch $Q$ of $C$ centered in (*$({\mathbb{C}}^{*})^{2}$, then we get ${\mathcal{M}}^{\prime}_{g,n}(\Delta,{\boldsymbol{w}})=\emptyset$. Thus, we are left with the following case: $\dim{\mathcal{M}}^{\prime}_{g,n}(\Delta)=n$, the curve $C$ is immersed in (*$({\mathbb{C}}^{*})^{2}$, and #$\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}=\#\Delta$. Arguing on the contrary, we show that $C$ is smooth along the toric divisors. Indeed, suppose that $C$ has a local singular branch $Q_{0}$ centered at some point $w\in{\operatorname{Tor}}(\sigma)$, $\sigma\in P^{1}_{\Delta}$. Fixing the position of the point $w$, we obtain a subfamily ${\mathcal{V}}_{1}\subset{\mathcal{M}}^{\prime}_{g,n}(\Delta)$ of dimension $\geq n-1$. On the other hand, the computations as above yield that
$\dim{\mathcal{V}}_{1}\leq h^{0}\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-D_{1})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{O}}_{{\operatorname{Tor}}(P_{\Delta})}(C)\right),$
where
$\deg D_{1}=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+\sum_{\sigma\in P^{1}_{\Delta}}\sum_{w\in C\cap{\operatorname{Tor}}(\sigma)}\sum_{Q\subset(C,w)}((Q\cdot{\operatorname{Tor}}(\sigma))_{w}-1)+{\operatorname{ord}}Q_{0},$
which implies
$\deg\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-D_{1})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{O}}_{{\operatorname{Tor}}(P_{\Delta})}(C)\right)=2g-2+\#\Delta-{\operatorname{ord}}Q_{0}\leq 2g-2+\#\Delta-2,$
and, by the Riemann-Roch formula, we obtain
$\dim{\mathcal{V}}_{1}\leq\#\Delta+g-1-2=n-2,$
which is a contradiction. Thus, $C$ must be totally immersed. Suppose now that some point $w\in C\cap{\operatorname{Tor}}(\sigma)$, $\sigma\in P^{1}_{\Delta}$, is a center of local branches $Q_{1},...,Q_{s}$, $s\geq 2$. Again, fixing the position of the point $w$, we obtain a subfamily ${\mathcal{V}}_{2}\subset{\mathcal{M}}^{\prime}_{g,n}(\Delta)$ of dimension $\geq n-1$. On the other hand, the computations as above yield that
$\dim{\mathcal{V}}_{2}\leq h^{0}\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-D_{2})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{O}}_{{\operatorname{Tor}}(P_{\Delta})}(C)\right),$
where
$\deg D_{2}=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+\sum_{\sigma\in P^{1}_{\Delta}}\sum_{w\in C\cap{\operatorname{Tor}}(\sigma)}\sum_{Q\subset(C,w)}((Q\cdot{\operatorname{Tor}}(\sigma))_{w}-1)+s,$
which implies
$\deg\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-D_{2})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{O}}_{{\operatorname{Tor}}(P_{\Delta})}(C)\right)=2g-2+\#\Delta-s\leq 2g-2+\#\Delta-2,$
and, by the Riemann-Roch formula, we obtain
$\dim{\mathcal{V}}_{2}\leq\#\Delta+g-1-2=n-2,$
which is a contradiction. Hence, $C$ is smooth along the toric divisors.
It remains to show that ${\mathcal{M}}_{g,n}(\Delta,{\boldsymbol{w}})$ does not contain elements $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]$ such that ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ is a multiple covering of the image $C^{\prime}={\Phi}(\widehat{C})$. Following the preceding argument, we assume that $\dim{\mathcal{M}}_{g,n}(\Delta)=n$ and that a generic element of some component ${\mathcal{V}}_{0}\subset{\mathcal{M}}_{g,n}(\Delta)$ is represented by an $s$-multiple covering ${\Phi}:\widehat{C}\to C^{\prime}\hookrightarrow{\operatorname{Tor}}(P_{\Delta})$, $s\geq 2$, and then come to a contradiction. Denote by ${\Phi}^{\prime}:\widehat{C}^{\prime}\to C^{\prime}$ the normalization and notice that the genus $g^{\prime}$ of $\widehat{C}^{\prime}$ is at most $g$, and that the number of local branches of $C^{\prime}$ centered on the toric divisors is bounded from above by #$\#\Delta$. Further on, we can assume that the variation of ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ along ${\mathcal{V}}_{0}$ induces an equisingular deformation of $C^{\prime}$ in the linear system $|C^{\prime}|$. The computations as in the preceding paragraph imply that
$n={\mathcal{V}}_{0}\leq\#\{Q\subset(C^{\prime},w),\ w\in C^{\prime}\cap\bigcup_{\sigma}{\operatorname{Tor}}(\sigma\}+g^{\prime}-1,$
which, in turn, yields that
#$\#\{Q\subset(C^{\prime},w),\ w\in C^{\prime}\cap\bigcup_{\sigma}{\operatorname{Tor}}(\sigma\}=\#\Delta\quad\text{and}\quad g^{\prime}=g.$
Moreover, the argument of the preceding paragraph would imply that $C^{\prime}$ is immersed and is smooth along the toric divisors. That means that ${\Phi}:\widehat{C}\to\widehat{C}^{\prime}$ is ramified at each of the #$\#\Delta\geq 3$ points of $\widehat{C}^{\prime}$ mapped to the toric divisors, and the ramification index at each of them equals $s$, but this contradicts the Riemann-Hurwitz formula. $\Box$
A map ${\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})$ such that $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{g,n}(\Delta,{\boldsymbol{w}})$ is said to be real if
  1. (i)
    the sequence ${\boldsymbol{w}}$ is invariant with respect to the tautological real structure on ${\operatorname{Tor}}(P_{\Delta})$,
  2. (ii)
    $(\widehat{C},{\boldsymbol{p}})$ is equipped with a real structure, and
  3. (iii)
    ${\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})$ commutes with the real structures in the source and in the target.
The set of equivalence classes of such real maps ${\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})$ considered up to equivariant isomorphism is denoted by ${\mathcal{M}}^{\mathbb{R}}_{g,n}(\Delta,{\boldsymbol{w}})$. We say that an element $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}^{\mathbb{R}}_{g,n}(\Delta,{\boldsymbol{w}})$ is separating if the complement in $\widehat{C}$ to the real points set ${\mathbb{R}}\widehat{C}$ is disconnected, i.e., consists of two connected components. The choice of one of these halves induces the so-called complex orientation on ${\mathbb{R}}\widehat{C}$, as well as on ${\Phi}({\mathbb{R}}\widehat{C})$ if ${\Phi}$ is birational onto its image. Denote by $\overrightarrow{\mathcal{M}}_{g,n}^{{\mathbb{R}}}(\Delta,{\boldsymbol{w}})$ the set of separating elements $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}^{\mathbb{R}}_{g,n}(\Delta,{\boldsymbol{w}})$ equipped with a choice of a half $\widehat{C}_{+}$ of $\widehat{C}\setminus{\mathbb{R}}\widehat{C}$.
Following [19], we assign to each element $\xi=([{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})],\widehat{C}_{+})\in\overrightarrow{\mathcal{M}}_{g,n}^{{\mathbb{R}}}(\Delta,{\boldsymbol{w}})$ its quantum index
${\operatorname{QI}}(\xi)=\frac{1}{\pi^{2}}\int_{\widehat{C}_{+}}{\Phi}^{*}\left(\frac{dx_{1}\wedge dx_{2}}{x_{1}x_{2}}\right),\quad x_{1}=|z_{1}|,\ x_{2}=|z_{2}|,$ (7)
with $z_{1},z_{2}$ coordinates in the torus (×$({\mathbb{C}}^{\times})^{2}\simeq{\operatorname{Tor}}(P_{\Delta})^{\times}$ such that the form $dx_{1}\wedge dx_{2}$ agrees with the orientation of ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ defined in Section 2.1. By [19, Theorem 3.1], if all intersection points of ${\Phi}(\widehat{C})$ with the toric divisors are real, then ${\operatorname{QI}}(\xi)\in\frac{1}{2}{\mathbb{Z}}$ and $|{\operatorname{QI}}(\xi)|\leq{\mathcal{A}}(\Delta)$.
3 Real refined enumerative invariants in genus 00, $1$, and $2$
3.1 Preliminary remarks
Recall that the count of real plane curves of a given degree and genus becomes non-invariant as long as the genus is positive [11, Theorem 3.1]. However, in the case of real rational surfaces with a disconnected real part, Welschinger invariants survive for certain positive genera $g$, provided that some conditions on constraints are fulfilled: in particular, each real connected component of the ambient surface contains at most one oval (i.e., global real branch) of any curve under consideration, and the counted curves have the maximal possible number $g+1$ of ovals; see [23]. We exploit here a similar approach in order to extend the invariant count of real rational curves [19] to positive genera. Our idea is to consider real curves that are tangent to toric divisors with even multiplicity at each intersection point. In such a case, every quadrant of a real toric surface plays the role of a separate component of the real part of a real surface. The lack of the “bad” wall in the space of constraints indicated in [11, Section 3] is expressed as a cohomology vanishing condition which we prove below in Lemmas 4.1, 5.1, and 6.1.
Remark 3.1
The above idea does not work for genus $g\geq 3$: if we fix constraints on the boundary of the positive quadrant (see Sections 3.2, 3.3, and 3.4) and $g$ more extra points in the big real torus, the set of counted curves will necessarily include non-dividing real curves with at most $g$ ovals, which does not allow one to perform Mikhalkin’s quantization.
3.2 Refined rational invariants
In the notation of Section 2.3, assume that
$g=0,\quad n_{in}=0,\quad n_{\partial}=n=\sum_{\sigma\in P^{1}_{\Delta}}n^{\sigma}-1,\quad\text{and}\quad{\boldsymbol{w}}={\boldsymbol{w}}_{\partial}\subset\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta}).$
In particular, for all $\sigma\in P^{1}_{\Delta}$ but one, $n(\sigma)=n^{\sigma}$, and for the remaining edge $\tau$, we have $n(\tau)=n^{\tau}-1$. Note that for a given ${\boldsymbol{w}}$, there exists a unique point $w^{\tau}_{n^{\tau}}\in\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ such that the sequence $\widehat{\boldsymbol{w}}:={\boldsymbol{w}}\cup\{w^{\tau}_{n^{\tau}}\}$ belongs to the $\operatorname{Men}(\Delta)$. We denote by ${\mathcal{M}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})\subset{\mathcal{M}}_{0,n}(\Delta,{\boldsymbol{w}})$ the subset formed by the curves passing through $\widehat{\boldsymbol{w}}$.
Denote by $\operatorname{Men}^{\prime}_{0}(\Delta)$ the subset of $\operatorname{Men}(\Delta)$ formed by the above sequences $\widehat{\boldsymbol{w}}$ such that ${\mathcal{M}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})$ satisfies the conclusions of Lemma 2.4. The closure
$\overline{\bigcup_{\widehat{\boldsymbol{w}}\in\operatorname{Men}_{0}(\Delta)}{\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})}$
is naturally fibered over the space of sequences $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)$, and for the sake of notation, we denote the fibers by ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})$. At last, we put
$\operatorname{Men}_{0}(\Delta)=\left\{\widehat{\boldsymbol{w}}\in\operatorname{Men}^{\prime}_{0}(\Delta)\ :\ \overline{\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})={\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})\right\}.$
Consider the subset ${\mathcal{M}}^{{\mathbb{R}},+}_{0,n}(\Delta,\widehat{\boldsymbol{w}})\subset{\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})$, formed by the elements $[{\Phi}:({\mathbb{P}}^{1},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]$ such that ${\Phi}({\mathbb{P}}_{\mathbb{R}}^{1})\subset{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ IIINote that ${\mathcal{M}}^{{\mathbb{R}},+}_{0,n}(\Delta,\widehat{\boldsymbol{w}})\subsetneq{\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})$ when there exists a connected component of ${\operatorname{Tor}}_{\mathbb{R}}(P_{\Delta})^{\times}\setminus{\operatorname{Tor}}_{\mathbb{R}}(P_{\Delta})^{+}$ such that the boundary of the component coincides with $\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$.. For every $[{\Phi}:({\mathbb{P}}^{1},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}^{{\mathbb{R}},+}_{0,n}(\Delta,\widehat{\boldsymbol{w}})$, the curve $C={\Phi}({\mathbb{P}}^{1})$ intersects each toric divisor ${\operatorname{Tor}}(\sigma)$ in points $w^{\sigma}_{i}$, $1\leq i\leq n^{\sigma}$. Hence, each element $\xi$ of the set
$\overrightarrow{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})=$$\{([{\Phi}:({\mathbb{P}}^{1},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})],{\mathbb{P}}^{1}_{+})\in\overrightarrow{\mathcal{M}}_{0,n}^{{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}})\ :$
$\qquad\qquad[{\Phi}:({\mathbb{P}}^{1},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})\}$
possesses a quantum index ${\operatorname{QI}}(\xi)\in\frac{1}{2}{\mathbb{Z}}$.
Furthermore, if $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{0}(\Delta)$, then (see [13, Section 1.1]), there is a well-defined Welschinger sign
[…]
where $e_{+}(C)$ is the number of elliptic nodes of $C={\Phi}({\mathbb{P}}^{1})$ in the positive quadrant ${\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$ that arise from a real nodal equigeneric deformation of all singular points of $C$ in ${\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$, and ${\varepsilon}(\xi,w_{i}^{\sigma})$ equals $1$ or $-1$ according as the complex orientation of ${\mathbb{R}}C={\Phi}({\mathbb{P}}_{\mathbb{R}}^{1})$ at $w_{i}^{\sigma}$ agrees or not with the fixed orientation of $\partial{\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$.
For each $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq{\mathcal{A}}(\Delta)$, put
$W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})=\sum_{\begin{array}[]{c}\scriptstyle{\xi\in\overrightarrow{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})}\\ \scriptstyle{{\operatorname{QI}}(\xi)=\kappa}\end{array}}W_{0}(\xi)\ .$
Theorem 3.2
For each $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq{\mathcal{A}}(\Delta)$, the value $W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ does not depend on the choice of a generic sequence $\widehat{\boldsymbol{w}}\subset\operatorname{Men}(\Delta)$.
We prove this theorem in Section 4.
Remark 3.3
Theorem 3.2 slightly generalizes the statement of [19, Theorem 5].
Remark 3.4
One can show ([19, 16]) that the invariants $W_{0}^{\kappa}(\Delta)$ vanish for all $\kappa\not\in 2{\mathbb{Z}}$ and for all $\kappa\in 2{\mathbb{Z}}$ such that $\kappa\not\equiv{\mathcal{A}}(\Delta)\mod 4$. In general, the invariants $W_{0}^{\kappa}(\Delta)$ can be computed via the formula in [2, Theorem 3.4].
3.3 Refined elliptic invariants
In the notation of Section 2.3, assume that
$g=1,\quad n_{in}=1,\quad n_{\partial}=\sum_{\sigma\in P^{1}_{\Delta}}n^{\sigma}-1,\quad{\boldsymbol{w}}_{\partial}\subset\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta}),\quad{\boldsymbol{w}}_{in}=\{w_{0}\}\subset{\mathcal{Q}},$
where ${\mathcal{Q}}$ is an open quadrant in ${\operatorname{Tor}}_{\mathbb{R}}(P_{\Delta})^{\times}\setminus{\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$. As in the preceding section, for all $\sigma\in P^{1}_{\Delta}$ but one, $n(\sigma)=n^{\sigma}$, and for the remaining edge $\tau$, we have $n(\tau)=n^{\tau}-1$. As in Section 3.2, we define the completion $\widehat{\boldsymbol{w}}_{\partial}:={\boldsymbol{w}}_{\partial}\cup\{w^{\tau}_{n^{\tau}}\}\in\operatorname{Men}(\Delta)$, where $w^{\tau}_{n^{\tau}}\in\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$, and put $\widehat{\boldsymbol{w}}=\widehat{\boldsymbol{w}}_{\partial}\cup{\boldsymbol{w}}_{in}$. We denote by ${\mathcal{M}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})\subset{\mathcal{M}}_{1,n}(\Delta,{\boldsymbol{w}})$ the subset formed by the curves passing through $\widehat{\boldsymbol{w}}$.
Denote by $\operatorname{Men}^{\prime}_{1}(\Delta)$ the set of sequences $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)\times{\mathcal{Q}}$ such that the set ${\mathcal{M}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})$ satisfies the conclusions of Lemma 2.4. The closure
$\overline{\bigcup_{\widehat{\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)}{\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})}$
is naturally fibered over the space of sequences $\widehat{\boldsymbol{w}}\subset\operatorname{Men}(\Delta)\times{\mathcal{Q}}$, and we denote the fibers by ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})$. At last, we introduce
$\operatorname{Men}_{1}(\Delta)=\left\{\widehat{\boldsymbol{w}}\in\operatorname{Men}^{\prime}_{1}(\Delta)\ :\ \overline{\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})={\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})\right\}.$
Consider the subset ${\mathcal{M}}^{{\mathbb{R}},+}_{1,n}(\Delta,\widehat{\boldsymbol{w}})\subset{\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})$, formed by the elements
$[{\Phi}:({\mathbb{E}},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})],\;\;\;\text{where}\;{\mathbb{E}}\;\text{is a smooth elliptic curve},$
defining real curves $C={\Phi}({\mathbb{E}})\subset{\operatorname{Tor}}(P_{\Delta})$ that have a one-dimensional real branch in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ and that intersect the toric divisor ${\operatorname{Tor}}(\sigma)$ in $w^{\sigma}_{i}$, $1\leq i\leq n^{\sigma}$, for all $\sigma\in P^{1}_{\Delta}$. It follows from Lemma 2.4 that for $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)$, the curve $C$ has two one-dimensional branches, one in ${\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$ and the other in the closed quadrant containing the point $w_{0}$. This defines a unique real structure on ${\mathbb{E}}$ whose fixed point set ${\mathbb{R}}{\mathbb{E}}$ consists of two ovals, making ${\Phi}:{\mathcal{E}}\to C$ a separating real curve. The choice of a component of ${\mathbb{E}}\setminus{\mathbb{R}}{\mathbb{E}}$ defines a complex orientation on the one-dimensional branches of ${\mathbb{R}}C$.
We impose a further restriction to the choice of the quadrant ${\mathcal{Q}}\supset\{w_{0}\}$ dictated by our wish to avoid specific wall-crossing events that may occur in the proof of the invariance of the count to be specified below: first, the escape of a real one-dimensional branch of counted curves out of the positive quadrant and, second, degenerations in which the union of the toric divisors splits off.
Definition 3.5
We say that the quadrant ${\mathcal{Q}}\subset{\operatorname{Tor}}_{\mathbb{R}}(P_{\Delta})^{\times}$ satisfies the admissible quadrant condition (AQC) if the closure $\overline{\mathcal{Q}}$ of ${\mathcal{Q}}$ shares with $\partial{\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$ at most one side.
Example 3.6
Let $d$, $d_{1}$ and $d_{2}$ be positive integers. Condition (AQC) imposes no restriction on the choice of ${\mathcal{Q}}$ in the case of the triangle ${\operatorname{conv}}\{(0,0),(d,0),(0,d)\}$ (the convex hull of the points $(0,0)$, $(d,0)$, $(0,d)$) having the projective plane ${\mathbb{P}}^{2}$ as associated toric surface, while for the rectangle ${\operatorname{conv}}\{(0,0),(d_{1},0),(d_{1},d_{2}),(0,d_{2})\}$ which has ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ as associated toric surface, the condition authorizes only the quadrant ${\mathcal{Q}}$ defined by the inequalities $x_{1}<0$ and $x_{2}<0$.
Denote by $\overrightarrow{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})$ the subset of $\overrightarrow{\mathcal{M}}_{1,n}^{{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}})$ formed by the elements
$([{\Phi}:({\mathbb{E}},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})],{\mathbb{E}}_{+})$
such that $[{\Phi}:({\mathbb{E}},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})$ and ${\mathbb{E}}_{+}$ is a half of ${\mathbb{E}}\setminus{\mathbb{R}}{\mathbb{E}}$. According to [19, Theorem 1], each element $\xi$ of $\overrightarrow{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})$ has a quantum index ${\operatorname{QI}}(\xi)\in\frac{1}{2}{\mathbb{Z}}$, $|{\operatorname{QI}}(\xi)|\leq{\mathcal{A}}(\Delta)$ (defined by formula (7) with the integration domain ${\mathbb{E}}_{+}$), and for ${\boldsymbol{w}}$ such that $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)$, according to the argument in [13, Section 1.1], there is a well-defined Welschinger sign
[…]
where $h(C,{\mathcal{Q}})$ is the number of hyperbolic nodes of $C={\Phi}({\mathbb{E}})$ in the quadrant ${\mathcal{Q}}$ that arise in any real nodal equigeneric deformation of all singularities of $C$ in ${\mathcal{Q}}$ (here $e_{0}(C)$ and ${\varepsilon}(\xi,w_{i}^{\sigma})$ are straightforward analogs of the corresponding ingredients in formula (8)).
For each $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq{\mathcal{A}}(\Delta)$, put
$W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})=\sum_{\begin{array}[]{c}\scriptstyle{\xi\in\overrightarrow{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})}\\ \scriptstyle{{\operatorname{QI}}(\xi)=\kappa}\end{array}}W_{1}(\xi)\ .$ (10)
Theorem 3.7
The value $W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ does not depend on the choice of a generic $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)\times{\mathcal{Q}}$ subject to the following conditions: $\widehat{\boldsymbol{w}}_{\partial}\subset\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ and $w_{0}\in{\mathcal{Q}}$, where ${\mathcal{Q}}$ is a fixed non-positive quadrant satisfying (AQC).
We prove this theorem in Section 5.
Remark 3.8
One can show [16] that the invariants $W_{1}^{\kappa}(\Delta,(\alpha,\beta))$ vanish for all $\kappa\not\in 2{\mathbb{Z}}$ and for all $\kappa\in 2{\mathbb{Z}}$ such that $\kappa\not\equiv{\mathcal{A}}(\Delta)\mod 4$.
3.4 Refined invariants of genus two
For genus $g=2$, we restrict our consideration to the case ${\operatorname{Tor}}(P_{\Delta})\simeq{\mathbb{P}}^{2}$. That is, $\Delta$ consists of positive multiples of the vectors $(1,1),(-1,0),(0,-1)$ and defines plane curves of an even degree $2d$. The boundary of the positive quadrant $({\mathbb{P}}^{2}_{\mathbb{R}})^{+}$ consists of the three segments $D_{s}^{+}$, $s=0,1,2$, where $D_{s}^{+}=D_{s}\cap{\operatorname{Cl}}({\mathbb{P}}_{\mathbb{R}}^{2})^{+}$, $D_{s}=\{x_{s}=0\}\subset{\mathbb{P}}^{2}$. We write ${\boldsymbol{w}}_{\partial}={\boldsymbol{w}}_{\partial}^{(1)}\cup{\boldsymbol{w}}_{\partial}^{(2)}\cup{\boldsymbol{w}}_{\partial}^{(0)}$, where
${\boldsymbol{w}}_{\partial}^{(1)}=(w_{i}^{(1)})_{1\leq i\leq n_{1}}\subset D_{1}^{+},\quad{\boldsymbol{w}}_{\partial}^{(2)}=(w_{i}^{(2)})_{1\leq i\leq n_{2}}\subset D_{2}^{+},\quad{\boldsymbol{w}}_{\partial}^{(0)}=(w_{i}^{(0)})_{1\leq i<n_{0}}\subset D_{0}^{+},$
and introduce the completions $\widehat{\boldsymbol{w}}_{\partial}={\boldsymbol{w}}_{\partial}\cup\{w^{(0)}_{n^{0}}\}\in\operatorname{Men}(\Delta)$ and $\widehat{\boldsymbol{w}}=\widehat{\boldsymbol{w}}_{\partial}\cup{\boldsymbol{w}}_{in}$, where ${\boldsymbol{w}}_{in}=\{w_{1},w_{2}\}$,
$w_{1}\in({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}=\left\{\frac{x_{1}}{x_{0}}<0,\ \frac{x_{2}}{x_{0}}>0\right\},\ w_{2}\in({\mathbb{P}}_{\mathbb{R}}^{2})^{\pm}=\left\{\frac{x_{1}}{x_{0}}>0,\ \frac{x_{2}}{x_{0}}<0\right\}.$
Denote by $\operatorname{Men}^{\prime}_{2}(\Delta)$ the set of sequences $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)\times({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}\times({\mathbb{P}}_{\mathbb{R}}^{2})^{\pm}$ such that the set ${\mathcal{M}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$ satisfies the conclusions of Lemma 2.4. The closure
$\overline{\bigcup_{\widehat{\boldsymbol{w}}\in\operatorname{Men}^{\prime}_{2}(\Delta)}{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})}$
is naturally fibered over the space of sequences $\widehat{\boldsymbol{w}}\subset\operatorname{Men}(\Delta)\times({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}\times({\mathbb{P}}_{\mathbb{R}}^{2})^{\pm}$, and we denote the fibers by ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$. Then we introduce
$\operatorname{Men}_{2}(\Delta)=\left\{{\boldsymbol{w}}\in\operatorname{Men}^{\prime}_{2}(\Delta)\ :\ \overline{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})={\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})\right\}.$
Denote by ${\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$ the set of real stable maps of curves of genus $2$ considered up to equivariant isomorphisms. Due to the choice of the points of $\widehat{\boldsymbol{w}}$, each element $[{\Phi}:(S_{2},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$, where $S_{2}$ stands for a real Riemann surface of genus $2$, defines an $M$-curve, i.e., ${\mathbb{R}}S_{2}$ consists of three ovals that are mapped to the quadrants ${\operatorname{Cl}}({\mathbb{P}}_{\mathbb{R}}^{2})^{+}$, ${\operatorname{Cl}}({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}$, ${\operatorname{Cl}}({\mathbb{P}}_{\mathbb{R}}^{2})^{\pm}$, respectively. The complement $S_{2}\setminus{\mathbb{R}}S_{2}$ consists of two components, and the choice of one of them, $S_{2}^{+}$ defined a complex orientation on ${\mathbb{R}}S_{2}$ and on ${\Phi}({\mathbb{R}}S_{2})$. Denote by $\overrightarrow{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$ the set of elements of ${\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,{\boldsymbol{w}})$ equipped with a complex orientation. Similarly to formula (9), we define for each element $\xi=[{\Phi}:(S_{2},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in\overrightarrow{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$,
$W_{2}(\xi)=(-1)^{e_{+}(C)}\cdot(-1)^{h_{-}(C)}\cdot\prod_{\begin{array}[]{c}\scriptstyle{k_{i}^{(s)}\equiv 0\mod 2}\\ \scriptstyle{1\leq i\leq n_{s},\ s=0,1,2}\end{array}}{\varepsilon}(\xi,w_{i}^{(s)}),$
where $C={\Phi}(S_{2})$, $e_{+}(C)$ is the number of elliptic nodes which appear in real local equigeneric nodal deformations of singular points of $C$ in $({\mathbb{P}}_{\mathbb{R}}^{2})^{+}$, $h_{-}(C)$ is the number of hyperbolic nodes which appear in real local equigeneric nodal deformations of singular points of $C$ in $({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}\cup({\mathbb{P}}_{\mathbb{R}}^{2})^{\pm}$, and ${\varepsilon}(\xi,w_{i}^{(s)})$ equals $1$ or $-1$ according as the complex orientation of the germ ${\Phi}(S_{2},p_{i}^{(s)})$ agrees with the positive orientation of $D_{s}^{+}$ or not. Recall that by [13, Lemma 2.2], the numbers $e_{+}(C)$ and $h_{-}(C)$ are well defined modulo $2$.
Theorem 3.9
For any toric degree $\Delta$ as above and any $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq 2d^{2}$, the number
$W^{\kappa}_{2}(\Delta,\widehat{\boldsymbol{w}})=\sum_{\begin{array}[]{c}\scriptstyle{\xi\in\overrightarrow{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})}\\ \scriptstyle{{\operatorname{QI}}(\xi)=\kappa}\end{array}}W_{2}(\xi)$
does not depend on the choice of a generic sequence $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)\times({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}\times({\mathbb{P}}_{\mathbb{R}}^{2})^{\pm}$.
We prove this theorem in Section 6.
4 Invariance in genus zero: proof of Theorem 3.2
Note that $\operatorname{Men}_{0}(\Delta)$ is an open dense semialgebraic subset of $\operatorname{Men}^{\tau}_{\mathbb{R}}(\Delta)$. We call its connected components the chambers. The complement to this subset is the union of finitely many semialgebraic strata of codimension $\geq 1$. We call the strata of codimension $1$ the walls.
Now we pick two sequences $\widehat{\boldsymbol{w}}(0)$ and $\widehat{\boldsymbol{w}}(1)$ satisfying the hypotheses of Theorem 3.2. We can join these sequences by a generic path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}\subset\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ which avoids the starta of codimension $\geq 2$. To prove Theorem 3.2, we need to verify the constancy of $W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ inside any given chamber and in all possible wall-crossing events, for all $\kappa$.
4.1 Invariance inside chambers
Here, we verify that $W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ does not change along intervals $t^{\prime}<t<t^{\prime\prime}$ lying in a given chamber. To this end, we show that the projection of ${\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}$ to the interval $(t^{\prime},t^{\prime\prime})$ does not have critical points. In such a case, the family ${\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}$ is the union of intervals trivially covering $(t^{\prime},t^{\prime\prime})$, and hence it lifts to a trivial family of complex oriented curves and the quantum index persists along each of the components of the latter family.
It is convenient to look at the elements $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ as follows: pick some pair $(\sigma_{0},i_{0})$, where $\sigma_{0}\in P^{1}_{\Delta}$ and $1\leq i_{0}\leq n^{\sigma_{0}}$, and think of $w^{\sigma_{0}}_{i_{0}}$ as a mobile point, whose position is determined by the other points $w^{\sigma}_{i}\in\widehat{\boldsymbol{w}}$ via the Menelaus condition. Put $C={\Phi}_{*}({\mathbb{P}}^{1})$. Observe that the tangent space at $C$ to the family of curves $C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|$ intersecting ${\operatorname{Tor}}(\sigma)$ at $w^{\sigma}_{i}$ (where $(\sigma,i)\neq(\sigma_{0},i_{0})$) with multiplicity at least $2k^{\sigma}_{i}$ can be identified with the linear system $\{C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|\ :\ (C^{\prime}\cdot C)_{w^{\sigma}_{i}}\geq 2k^{\sigma}_{i}\}$. In turn, the tangent space at $C$ to the family of curves $C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|$ intersecting ${\operatorname{Tor}}(\sigma_{0})$ at a point close to $w^{\sigma_{0}}_{i_{0}}$ with multiplicity at least $2k^{\sigma_{0}}_{i_{0}}$ can be identified with the linear system $\{C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|\ :\ (C^{\prime}\cdot C)_{w^{\sigma_{0}}_{i_{0}}}\geq 2k^{\sigma}_{i}-1\}$. Then, the required transversality amounts to the following statement.
Lemma 4.1
Let $\xi=[{\Phi}:({\mathbb{P}}^{1},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})$ with $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{0}(\Delta)$. Then,
$H^{0}\left({\mathbb{P}}^{1},{\mathcal{N}}_{\Phi}\left(-\sum_{\sigma\neq\sigma_{0},i\neq i_{0}}2k^{\sigma}_{i}p^{\sigma}_{i}-(2k^{\sigma_{0}}_{i_{0}}-1)p^{\sigma_{0}}_{i_{0}}\right)\right)$
$=H^{1}\left({\mathbb{P}}^{1},{\mathcal{N}}_{\Phi}\left(-\sum_{\sigma\neq\sigma_{0},i\neq i_{0}}2k^{\sigma}_{i}p^{\sigma}_{i}-(2k^{\sigma_{0}}_{i_{0}}-1)p^{\sigma_{0}}_{i_{0}}\right)\right)=0.$
Proof. We have
[…]
Thus, the claim of the lemma follows from the Riemann-Roch theorem. $\Box$
4.2 Walls in the space of constraints
Here, we characterize elements $\xi\in\overline{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{0\leq t\leq 1}$ that are generic elements of strata of dimension $n-1$ (i.e., of codimension $1$) in $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)\setminus\operatorname{Men}_{0}(\Delta)$.
Lemma 4.2
(1) The following elements $\xi=[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]$ cannot occur in ¯$\overline{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{0\leq t\leq 1}\setminus{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{0\leq t\leq 1}$:
  1. (1i)
    $\widehat{C}$ is a reducible connected curve of arithmetic genus 00 with a component mapped onto a toric divisor;
  2. (1ii)
    $\widehat{C}$ is a reducible connected curve of arithmetic genus 00 with at least three irreducible components.
(2) If $\xi=[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in\overline{\mathcal{M}}_{0,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$, where $\widehat{\boldsymbol{w}}(t^{*})$ is a generic element in an $(n-1)$-dimensional stratum in $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)\setminus\operatorname{Men}_{0}(\Delta)$, then $\xi$ is of one of the following types:
  1. (2i)
    $\widehat{\boldsymbol{w}}(t^{*})$ consists of $n+1$ distinct points, $\widehat{C}\simeq{\mathbb{P}}^{1}$, the map ${\Phi}$ is birational onto its image and satisfies the following: either it is smooth at $\widehat{\boldsymbol{w}}(t^{*})$, but has singular branches in ${\operatorname{Tor}}(P_{\Delta})^{\times}$ or it is an immersion everywhere but at one point $w_{i}^{\sigma}(t^{*})$, where it has a singularity of type $A_{2m}$, $m\geq k_{i}^{\sigma}$, and, furthermore, $C={\Phi}(\widehat{C})$ is smooth at each point of $\widehat{\boldsymbol{w}}(t^{*})\setminus\{w_{i}^{\sigma}(t^{*})\}$;
  2. (2ii)
    two points of the sequence $\widehat{\boldsymbol{w}}(t^{*})$ coincide ($w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})$ for some $\sigma\in P^{1}_{\Delta}$ and $i\neq j$) and $\widehat{C}\simeq{\mathbb{P}}^{1}$, the map ${\Phi}$ being an immersion such that the point $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})$ is a center of one or two smooth branches;
  3. (2iii)
    $\widehat{\boldsymbol{w}}(t^{*})$ consists of $n+1$ distinct points, $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $\widehat{C}_{1}\simeq\widehat{C}_{2}\simeq{\mathbb{P}}^{1}$ and $\widehat{C}_{1}\cap\widehat{C}_{2}$ is one point $p$, each map ${\Phi}:\widehat{C}_{i}\to{\operatorname{Tor}}(P_{\Delta})$, $i=1,2$, is either an immersion smooth along ${\operatorname{Tor}}(\partial P_{\Delta})$, or a multiple covering of a line intersecting only two toric divisors, while these divisors correspond to opposite parallel sides of $P_{\Delta}$ and the intersection points with these divisors are ramification points of the covering; furthermore, either ${\Phi}(p)\in{\operatorname{Tor}}(P_{\Delta})^{\times}$ and then all intersection points of the curves $C_{1}={\Phi}(\widehat{C}_{1})$, $C_{2}={\Phi}(\widehat{C}_{2})$ are ordinary nodes, or $w={\Phi}(p)\in{\boldsymbol{w}}(t^{*})$, and in the latter case at least one of the maps ${\Phi}:\widehat{C}_{i}\to{\operatorname{Tor}}(P_{\Delta})$ is birational onto its image, the curves $C_{1}={\Phi}(\widehat{C}_{1})$, $C_{2}={\Phi}(\widehat{C}_{2})$ intersect ${\operatorname{Tor}}(\sigma)$ at $w$ with even multiplicities, and they do not have common point in $\widehat{\boldsymbol{w}}(t^{*})\setminus\{w\}$;
  4. (2iv)
    $P_{\Delta}$ is a triangle, $n=3$, two points of the sequence $\widehat{\boldsymbol{w}}(t^{*})$ coincide, the curve $C={\Phi}({\mathbb{P}}^{1})$ is rational and smooth at $\widehat{\boldsymbol{w}}(t^{*})$, and ${\Phi}:{\mathbb{P}}^{1}\to C$ is a double covering ramified at two distinct points of $\widehat{\boldsymbol{w}}(t^{*})$.
Proof. (1i) A toric divisor ${\operatorname{Tor}}(\sigma)$ cannot split off alone, since, otherwise, in the deformation along the path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$, the intersection points with the neighboring toric divisors would yield points on these toric divisors on the distance less than $\tau$ from the corners of ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$. For the same reason we obtain that only all toric divisors together may split off, while their intersection points must smooth out in the deformation. However, this contradicts the rationality of the considered curves.
(1ii) First, note that ${\Phi}(\widehat{C})$ intersects ${\operatorname{Tor}}(\partial P_{\Delta})$ only at $\widehat{\boldsymbol{w}}(t^{*})$. Furthermore, if the images of two irreducible components $\widehat{C}_{1},\widehat{C}_{2}$ of $\widehat{C}$ contain the same point $w_{i}^{\sigma}(t^{*})$ that is different from any other point $w_{j}^{\sigma}(t^{*})$, $j\neq i$, then $w_{i}^{\sigma}(t^{*})={\Phi}(\widehat{C}_{1}\cap\widehat{C}_{2})$ (cf. Lemma 2.4). Due to the genus restriction, it follows that the sequence $\widehat{\boldsymbol{w}}(t^{*})$ satisfies at least three Menelaus conditions, which cuts off $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ a polytope of dimension $\leq n-2$, contrary to the dimension $n-1$ assumption.
Let $\xi$ satisfy the hypotheses of item (2).
(2i) Let $\widehat{C}\simeq{\mathbb{P}}^{1}$ and $\widehat{\boldsymbol{w}}(t^{*})$ consist of $n+1$ distinct points. If all points $w_{i}^{\sigma}(t^{*})$ are smooth, then we get the case (2i).
Suppose that $C={\Phi}({\mathbb{P}}^{1})$ is singular at some point $w_{i}^{\sigma}(t^{*})$. Note that $\xi$ must be unibranch at each point of $\widehat{\boldsymbol{w}}(t^{*})$. Since $\xi$ is a genetic element of an $(n-1)$-dimensional family in ${\mathcal{M}}_{0,n}(\Delta)$, fixing the position of $w_{i}^{\sigma}$, we obtain a family of dimension $\geq n-2$. Applying [13, Inequality (5) in Lemma 2.1], we obtain
$\sum_{\sigma\subset\partial P}\|\sigma\|_{\mathbb{Z}}\geq 2+\left(\sum_{\sigma\subset\partial P}\|\sigma\|_{\mathbb{Z}}-n\right)+(n-3)+\sum_{B}({\operatorname{ord}}B-1)$
$=\sum_{\sigma\subset\partial P}\|\sigma\|_{\mathbb{Z}}-1+\sum_{B}({\operatorname{ord}}B-1)\ ,$
where $B$ ranges over all singular branches of $\xi$, and hence $\xi$ has a unique singular branch, and this branch is centered at $w_{i}^{\sigma}(t^{*})$ and has multiplicity $2$. Thus, the type of the singularity must be $A_{2m}$, $m\geq k_{i}^{\sigma}$.
(2ii) Suppose that $\widehat{C}\simeq{\mathbb{P}}^{1}$ and some of the points of the sequence $\widehat{\boldsymbol{w}}(t^{*})$ coincide. For the dimension reason, the number #$\#(\widehat{\boldsymbol{w}}(t^{*}))$ of points in $\widehat{\boldsymbol{w}}(t^{*})$ is equal to $n-1$ (i.e., $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})=w$ for some $\sigma\subset\partial P$ and $i\neq j$), and all these $n-1\geq 2$ points are in general position subject to the Menelaus relation (1). If $n-1=2$, then the claim of item (2ii) is fulfilled by [22, Lemma 3.5]. Assume that $n-1\geq 3$. If $C={\Phi}(\widehat{C})$ is unibranch at each point of $\widehat{\boldsymbol{w}}(t^{*})$, then, by Lemma 2.4, the curve $C$ is immersed and smooth along the toric divisors. If $C$ is not unibranch at $w$, then by Lemma 2.4 it has two local branches at $w$; furthermore, $C$ must be unibranch at each point $w_{i^{\prime}}^{\sigma^{\prime}}(t^{*})\neq w$ and, in addition, smooth if $k_{i^{\prime}}^{\sigma^{\prime}}\geq 2$ due to Lemma 2.4 and claim (1ii). Let us show that $C$ is immersed. Fixing the position of $w$ and the position of one more point $w^{\prime}\in{\boldsymbol{w}}(t^{*})\setminus\{w\}$, we obtain a family of dimension at least $n-3\geq 1$; thus, [13, Inequality (5) in Lemma 2.1] applies:
$c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})\geq 2+(c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})-n+2)+\sum_{B}({\operatorname{ord}}B-1)+(n-4)$
$=c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})+\sum_{B}({\operatorname{ord}}B-1)\ ,$
where $B$ runs over all singular local branches of $C$ in ${\operatorname{Tor}}(P_{\Delta})^{\times}\cup\{w,w^{\prime}\}$, and hence $C$ is immersed.
(2iii) We are left with the case of $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $\widehat{C}_{1}\simeq\widehat{C}_{2}\simeq{\mathbb{P}}^{1}$ and $\widehat{C}_{1}\cap\widehat{C}_{2}=\{p\}$ is one point. For the dimension reason, the points of $\widehat{\boldsymbol{w}}(t^{*})$ are in general position subject to exactly two Menelaus conditions (induced by the components of $\widehat{C}$), and they are all distinct. Each of the curves $C_{1}$, $C_{2}$, which passes through at least three points of $\widehat{\boldsymbol{w}}(t^{*})$, is immersed and smooth along the toric divisors and has even intersection multiplicity with ${\operatorname{Tor}}(\partial P_{\Delta})$ at any point, by Lemma 2.4. If $C_{1}$ or $C_{2}$ passes through exactly two points of $\widehat{\boldsymbol{w}}(t^{*})$, then it is a multiple covering of a line as described in item (2iv). Let ${\Phi}(p)\in{\operatorname{Tor}}(P_{\Delta})^{\times}$. Then, by Lemma 2.4, the curves $C_{1}$ and $C_{2}$ do not share points in $\widehat{\boldsymbol{w}}(t^{*})$, and we claim that all their intersection points are ordinary nodes. Due to the genericity assumptions for $\widehat{\boldsymbol{w}}(t^{*})$, we have to study the only case of both $C_{1}$ and $C_{2}$ immersed. If we freely move the points of $C_{1}\cap\widehat{\boldsymbol{w}}(t^{*})$ so that the corresponding Menelaus condition induced by $C_{1}$ retains, and fix the curve $C_{2}$, then the persisting tangency condition of a germ ${\Phi}:(\widehat{C}_{1},q)\to{\operatorname{Tor}}(P_{\Delta})$ to the curve $C_{2}$ would yield that the tangent space to the considered family of curves in the linear system $|C_{1}|$ would be contained in $H^{0}(\widehat{C}_{1},{\mathcal{O}}_{\widehat{C}_{1}}({\boldsymbol{d}}_{1}))$, where
$\deg{\boldsymbol{d}}_{1}=C_{1}^{2}-(C_{1}^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta}))[C_{1}]+2)-(c_{1}({\operatorname{Tor}}(P_{\Delta}))[C_{1}]-\#({\boldsymbol{w}}(t^{*})\cap C_{1}))-1$
=#$=\#(\widehat{\boldsymbol{w}}(t^{*})\cap C_{1})-3>-2\ ,$
and hence, by the Riemann-Roch theorem,
$h^{0}(\widehat{C}_{1},{\mathcal{O}}_{\widehat{C}_{1}}({\boldsymbol{d}}_{1}))=\#(\widehat{\boldsymbol{w}}(t^{*})\cap C_{1})-3+1=\#(\widehat{\boldsymbol{w}}(t^{*})\cap C_{1})-2<\#\widehat{\boldsymbol{w}}(t^{*})\cap C_{1})-1\ ,$
which is a contradiction. If ${\Phi}(p)=w_{i}^{\sigma}(t^{*})\in\widehat{\boldsymbol{w}}(t^{*})$, then the intersection multiplicities of the germs ${\Phi}(\widehat{C}_{1},p)$ and ${\Phi}(\widehat{C}_{2},p)$ with ${\operatorname{Tor}}(\sigma)$ are even. Indeed, otherwise, $C_{1}$ and $C_{2}$ would have at least one additional local branch centered on some ${\operatorname{Tor}}(\sigma^{\prime})$, $\sigma^{\prime}\neq\sigma$, and intersecting ${\operatorname{Tor}}(\sigma^{\prime})$ with an odd multiplicity, which contradicts our assumptions.
(2iv) Suppose that ${\Phi}:{\mathbb{P}}^{1}\to{\operatorname{Tor}}(P_{\Delta})$ is an $s$-multiple covering onto its image, $s\geq 2$. For the dimension reason, the sequence $\widehat{\boldsymbol{w}}(t^{*})$ contains at least $n$ distinct points. If $\widehat{\boldsymbol{w}}(t^{*})$ consists of $n-1$ distinct points, then at each point of $\widehat{\boldsymbol{w}}(t^{*})$ we have the ramification index $s$; hence, by the Riemann-Hurwitz formula, we get
$2\leq 2s-(s-1)(n+1),$ (11)
which gives $n\leq 1$; the latter inequality contradicts the fact that $n+1$ is bounded from below by the number of sides of $P_{\Delta}$. If the sequence $\widehat{\boldsymbol{w}}(t^{*})$ contains only $n$ distinct points, we have $n\geq 3$. Note that ${\Phi}$ has an irreducible preimage in at least $n-1$ points of $\widehat{\boldsymbol{w}}(t^{*})$ with ramification index $s$, and hence $n=3$ (cf. (11)). It also follows that there are no other ramifications and that $s=2$, since the remaining point of $\widehat{\boldsymbol{w}}(t^{*})$, where ${\Phi}$ is not ramified, lifts to at most two points in ${\mathbb{P}}^{1}$. Thus, we are left with the case described in item (2iv) of the lemma. $\Box$
4.3 Wall-crossing
We complete the proof of Theorem 3.2 with the following lemma.
Lemma 4.3
Let $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ be a generic path in $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$, and let $t^{*}\in(0,1)$ be such that ¯$\overline{\mathcal{M}}_{0,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ contains an element $\xi$ as described in one of the items of Lemma 4.2(2). Then, for each $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq{\mathcal{A}}(\Delta)$, the numbers $W_{0}^{\kappa}(t)=W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ do not change as $t$ varies in a neighborhood of $t^{*}$.
Proof. We always can assume that in a neighborhood of $t^{*}$, the path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ is defined by fixing the position of some $n-1$ points of $\widehat{\boldsymbol{w}}(t^{*})$, while the other two points remain mobile (the choice of the two mobile points may depend on the considered degeneration). We also notice that, in the degenerations as in Lemma 4.2(2i,2ii), the source curve and its real structure remain fixed, which implies that the quantum index is constant in these wall-crossings. Except for the case of Lemma 4.2(2iii) describing reducible degenerations, we work with families of curves which are trivially covered by families of complex oriented curves so that the quantum index persists along each component of the family of oriented curves.
(1) Suppose that $\xi\in\overline{\mathcal{M}}^{{\mathbb{R}},+}_{0,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ is as in Lemma 4.2(2i) and, moreover, if it has a singular branch at $w_{i}^{\sigma}(t^{*})$, then $k_{i}^{\sigma}=1$. We derive the constancy of $W_{0}^{\kappa}(t)$, $|t-t^{*}|<\delta$ from [12, Lemma 15] and [13, Lemma 2.4(1)]. To this end, we have to establish the following transversality statement (cf. [12, Lemma 13]). Choose a sufficiently large integer $s$. For each point $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}$, we set $I_{z}=I^{cond}(C,z)/{\mathfrak{m}}_{z}^{s}\subset{\mathcal{O}}_{C,z}/{\mathfrak{m}}_{z}^{s}$, the quotient of the conductor ideal by the power of the maximal ideal, which can be viewed as the tangent cone to the stratum parameterizing equigeneric deformations (see [3, Theorem 4.15]). For each point $w=w_{i}^{\sigma}(t^{*})$, where $C$ is smooth and which is fixed, respectively, mobile, we set
$I_{w}=\{\varphi\in{\mathcal{O}}_{C,w}\ :\ (\varphi\cdot C)_{w}\geq 2k_{i}^{\sigma}\}/{\mathfrak{m}}_{w}^{s}\subset{\mathcal{O}}_{C,w}/{\mathfrak{m}}_{w}^{s}\ ,$
[…]
which can be viewed as the tangent space to the stratum parameterizing deformations that keep the intersection number with ${\operatorname{Tor}}(\sigma)$ at $w$, respectively, in a nearby point on ${\operatorname{Tor}}(\sigma)$. For each point $w=w_{i}^{\sigma}(t^{*})$ with $k_{i}^{\sigma}=1$, where $C$ is singular (i.e., has a singular branch of type $A_{2m}$, $m\geq 1$) and which is fixed, respectively, mobile, we set
$I_{w}=\{\varphi\in{\mathcal{O}}_{C,w}\ :\ (\varphi\cdot C)_{w}\geq 2+2m\}/{\mathfrak{m}}_{w}^{s}\subset{\mathcal{O}}_{C,w}/{\mathfrak{m}}_{w}^{s}\ ,$
[…]
which can be viewed as the tangent cone to the stratum parameterizing equigeneric deformations with the fixed intersection multiplicity $2$ with ${\operatorname{Tor}}(\sigma)$ at $w$, respectively, at a nearby point on ${\operatorname{Tor}}(\sigma)$ (see [13, Lemma 2.4(1)] and [24, Lemma 3(1)]). The required transversality is as follows: the natural image of the germ of the linear system $|{\mathcal{L}}_{P_{\Delta}}|$ at $C$ to $\prod_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}{\mathcal{O}}_{C,z}/{\mathfrak{m}}_{z}^{s}\times\prod_{w\in{\boldsymbol{w}}(t^{*})}{\mathcal{O}}_{C,w}/{\mathfrak{m}}_{w}^{s}$ intersects there transversally with $\prod_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}I_{z}\times\prod_{w\in\widehat{\boldsymbol{w}}(t^{*})}I_{w}$. The cohomological reformulation amounts to the relation
$H^{0}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}({\boldsymbol{d}}))=0,\quad{\boldsymbol{d}}\in{\operatorname{Div}}({\mathbb{P}}^{1})\ ,$ (12)
where the points of ${\boldsymbol{w}}(t^{*})$ are fixed and the remaining point of $\widehat{\boldsymbol{w}}(t^{*})$ is mobile. Since
$\deg{\boldsymbol{d}}=C^{2}-\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}2\delta(C,z)-\sum_{w\in\widehat{\boldsymbol{w}}(t^{*})}\dim{\mathcal{O}}_{C,w}/I_{w}$
$=C^{2}-(C^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})+2)-(c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})-1)=-1>-2\ ,$
and hence (12) by the Riemann-Roch formula. The established transversality reduces the constancy of $W_{0}^{\kappa}(t)$ to the constancy of the count of Welschinger signs when varying germs $(C,z)$ and $(C,w)$ for $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}$ and $w\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap\widehat{\boldsymbol{w}}(t^{*})$. Thus, the constancy in the former variation follows from [12, Lemma 15] (see also [12, Lemma 13]), while the constancy in the latter variation follows from the fact that the equigeneric stratum, which we consider in ${\mathcal{O}}_{C,w}/{\mathfrak{m}}_{w}^{s}$, is smooth (see [24, Lemma 3(1)] and [13, Lemma 2.4(2)]), and the local count of Welschinger signs is invariant by [13, Lemma 2.4(1)].
(2) Suppose that $\xi\in\overline{\mathcal{M}}^{{\mathbb{R}},+}_{0,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ is as in Lemma 4.2(2i) with a singularity at a point $w_{i}^{\sigma}(t^{*})$ of type $A_{2m}$, $m\geq k_{i}^{\sigma}$. We can assume that the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<{\varepsilon}}$ is such that $n-1$ points of $\widehat{\boldsymbol{w}}(t)$, including $w_{i}^{\sigma}(t)$, stay fixed, whereas the other two points move keeping the Menelaus condition. We claim that the germ at $\xi$ of the family {¯$\{\overline{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))\}_{|t-t^{*}|<{\varepsilon}}$ is smooth and regularly parameterized by $t\in(t^{*}-{\varepsilon},t^{*}+{\varepsilon})$. By [3, Theorem 4.15 and Proposition 4.17(2)] and [24, Lemma 3(1)], it is sufficient to prove that
$H^{1}(C,{\mathcal{J}}_{Z/C}(C))=0\ ,$ (13)
where $C={\Phi}({\mathbb{P}}^{1})$ and ${\mathcal{J}}_{Z/C}\subset{\mathcal{O}}_{C}$ is the ideal sheaf of the zero-dimensional scheme $Z\subset C$ concentrated at $\widehat{\boldsymbol{w}}(t^{*})\cup{\operatorname{Sing}}\hskip 1.70709pt(C)$ and given (cf. Step (1) of the proof)
  • by the ideal $I_{w}=\{\varphi\in{\mathcal{O}}_{C,w}\ :\ {\operatorname{ord}}\varphi\big{|}_{w}\geq 2k_{i^{\prime}}^{\sigma^{\prime}}\}$ at each point $w=w_{i^{\prime}}^{\sigma^{\prime}}(t^{*})\in\widehat{\boldsymbol{w}}(t^{*})\setminus\{w_{i}^{\sigma}(t^{*})\}$, which is fixed as $|t-t^{*}|<{\varepsilon}$,
  • by the ideal $I_{w}=\{\varphi\in{\mathcal{O}}_{C,w}\ :\ {\operatorname{ord}}\varphi\big{|}_{w}\geq 2k_{i^{\prime}}^{\sigma^{\prime}}-1\}$ at each point $w=w_{i^{\prime}}^{\sigma^{\prime}}(t^{*})\in\widehat{\boldsymbol{w}}(t^{*})\setminus\{w_{i}^{\sigma}(t^{*})\}$, which moves as $|t-t^{*}|<{\varepsilon}$,
  • by the ideal $I_{w}=\{\varphi\in{\mathcal{O}}_{C,w}\ :\ {\operatorname{ord}}\varphi\big{|}_{w}\geq 2k_{i}^{\sigma}+2m\}$ at the point $w=w_{i}^{\sigma}(t^{*})$ (here, in the notations of [24, Lemma 3(1)], $s=2k_{i}^{\sigma}$ and $\delta(C,w)=m$),
  • by the conductor ideal $I^{cond}_{C,z}$ at each point $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\setminus\{w_{i}^{\sigma}(t^{*})\}$ (for the definition of the conductor ideal see, for instance, [3, Section 1, item (iii)] and [9, Section I.3.4, item “Semigroup and Conductor”]).
Thus, the desired relation (13) turns into
$H^{1}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}({\boldsymbol{d}}))=0\ ,$
(cf. (12)), where the divisor ${\boldsymbol{d}}$ has degree
$\deg{\boldsymbol{d}}=C^{2}-\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\setminus{\boldsymbol{w}}(t^{*})}\dim{\mathcal{O}}_{C,z}/I_{z}-\sum_{w\in{\boldsymbol{w}}(t^{*})}\dim{\mathcal{O}}_{C,w}/I_{w}$
$=C^{2}-(C^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})+2)-(c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})-2)=0>-2\ ,$
which, finally, confirms (13).
[…]
$F_{t_{1}}(x,y):=F(x,y)+\sum_{j=0}^{k-1}t_{1}^{(k-j)\gamma}(a_{j1}+O(t_{1}))x^{j}y+\sum_{p+kq\geq 2k}b_{pq}(t_{1})x^{p}y^{q}=0\ ,$ (14)
where $a_{01}\neq 0$ and $b_{pq}(0)=0$ for all $p,q$ in the range. The tropical limit of the germs $[{\Phi}_{t}({\mathbb{P}}^{1}),w]$ as $t\to t^{*}$ or, equivalently, $t_{1}\to 0$ defines the subdivision of the Newton polygon $P(F_{t})$ of $F_{t}(x,y)$ into the triangle $T_{1}={\operatorname{conv}}\{(0,1),(0,2),(2k,0)\}$ and the Newton polygon $P(F)$ of $F$, see Figure 1(a) (the details about the tropicalization can be found in [18, Section 3] and [15, Section 2.5.4], see also [22, Section 2] and [5, Section 2.2]). Note that, since the curves ${\Phi}_{t}({\mathbb{P}}^{1})$ are rational, the triangle $T_{1}$ is not subdivided further, and the corresponding limit curve
$C_{1}:=\left\{\sum_{j=0}^{k-1}a_{j1}x^{j}y+y^{2}-2x^{k}y+x^{2k}=0\right\}$ (15)
is rational. Moreover, since the curve $C$ is unibranch at $w$, the limit curve $C_{F}$ defined by the polynomial $F$ in the surface ${\operatorname{Tor}}(P(F))$ is unibranch at the intersection point with the toric divisor ${\operatorname{Tor}}([(0,2),(2k,0)])$. Hence, $C_{1}$ is unibranch at its intersection point with the toric divisor ${\operatorname{Tor}}([(0,2),(2k,0)])$, because otherwise, the union of the limit curves $C_{1}$ and $C_{F}$ would deform into a non-rational curve contrary to the fact that the curves ${\Phi}({\mathbb{P}}^{1})$ are rational. By [22, Lemma 3.5], the curve $C_{1}$ is smooth at the intersection points with the toric divisors.
[…]
Figure 1: Proof of Lemma 4.3, part (2), I
We perform the modification of the above tropicalization along the toric divisor ${\operatorname{Tor}}([(0,2),(2k,0)])$, in the sense of [20, Section 5] (see also [5, Section 2.2]). Recall that the curve $C$ has singularity of type $A_{2m}$ at $w$, and hence there exists a coordinate change
$x=x^{\prime},\quad y=y^{\prime}+(x^{\prime})^{k}+\sum_{k<j\leq m}c_{j}(x^{\prime})^{j}\ ,$ (16)
which converts the polynomial $F(x,y)$ into a polynomial
$F^{\prime}(x^{\prime},y^{\prime})=(y^{\prime})^{2}+a_{2m+1,0}(x^{\prime})^{2m+1}+\text{h.o.t.},\quad a_{2m+1,0}\neq 0\ .$
There exists a deformation of the coordinate change (16)
$x=x^{\prime},\quad y=y^{\prime}+(1+O(t_{1}))x^{k}+\sum_{k<j<m}(c_{j}+O(t_{1}))x^{j}\ ,$ (17)
which turns the family of polynomials $F_{t_{1}}(x,y)$, $|t_{1}|<{\varepsilon}$, into the family
$F^{\prime}_{t_{1}}(x^{\prime},y^{\prime})=\sum_{j=0}^{k-1}t_{1}^{(k-j)a}(a_{j1}+O(t_{1}))(x^{\prime})^{j}y^{\prime}+\sum_{j=2k-1}^{2m}t_{1}^{\nu(j)}(a_{j0}+O(t_{1}))(x^{\prime})^{j}$
[…]
The tropical limit of the latter family of polynomials defines the subdivision of the Newton polygon $P(F\_{t_{1}})$ into the triangles
$T_{1}={\operatorname{conv}}\{((0,2),(0,1),(2k-1,0)\},\quad T_{3}={\operatorname{conv}}\{(0,2),(2k-1,0),(2m+1,0)\}\ ,$
and the Newton polygon $P(F^{\prime})$ (see Figure 1(b)). Moreover, in principle, the triangle $T_{3}$ can be subdivided further. Taking into account that all limit curves associated with the considered tropical limit must be rational and that the coefficients of $F^{\prime}_{t_{1}}$ at the interior integral points of $T_{3}$ vanish, the triangle $T_{3}$ can be subdivided only into triangles of type ${\operatorname{conv}}\{(0,2),(2j^{\prime}-1,0),(2j^{\prime\prime}+1,0)\}$ with some $k\leq j^{\prime}\leq j^{\prime\prime}\leq m$. The function $\nu:\{2k-1,2k,...,2m+1\}\to{\mathbb{Z}}$ in the exponents of $t_{1}$ in (18) (with the extra value $\nu(2m+1)=0$) extends to a convex piecewise-linear function on the whole segment $[(2k-1,0),(2m+1,0)]$ with linearity domains induced by the subdivision of the triangle $T_{3}$. We make two remarks.
  • Since the germ at $\xi$ of the family {¯$\{\overline{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))\}_{|t-t^{*}|<{\varepsilon}}$ is smooth and regularly parameterized by $t_{1}$ (see the beginning of Step (2)), we derive that $\mu(2m)=1$.
  • Since the linearity domains of $\nu$ are segments of even length, the value $\nu(2k-1)$ is even. On the other hand, formulas (17) yield that $\nu(2k-1)=\alpha$, and hence all exponents of $t_{1}$ in the leading terms of the coefficients of $F_{t_{1}}$ at $x^{j}y$, $10\leq j\leq 2k-$ 1, are even (see (14)). This means, in particular, that the limit curve $C_{1}$ given by (15), contributes to the curves ${\Phi}_{t}({\mathbb{P}}^{1})$ the same collection of real nodes with the same location with respect to ${\operatorname{Tor}}(\sigma)$ both for $t<t^{*}$ and for $t>t^{*}$. Hence, these singularities contribute the same factor to $W_{0}(\xi_{t})$ for $t<t^{*}$ and for $t>t^{*}$ (cf. formula (8)).
Now we analyze the contribution of the limit curves associated with the fragment $T_{3}$, and we consider the only case when $T_{3}$ is not subdivided. The other cases can be treated similarly. So, we have the limit curve
$C_{3}=\left\{(y^{\prime})^{2}+\sum_{j=2k-1}^{2m+1}a_{j0}(x^{\prime})^{j}=0\right\}\ ,$
which must be rational, and this yields that
$\sum_{j=2k-1}^{2m+1}a_{j0}(x^{\prime})^{j}=a_{2m+1,0}(x^{\prime})^{2k-1}Q(x^{\prime})^{2},\quad\deg Q=r:=m-k+1\ ,$
while the nodes of $C_{3}$ are $(\lambda_{j},0)$, $j=1,...,r$, with the numbers $\lambda_{j}$ being the roots of $Q$. Denote by $r_{+}$ (respectively, $r_{-}$) the number of positive (respectively, negative) roots of $Q$. Without loss of generality we can assume that $a_{2m+1,0}<0$. Then, the curve $C_{3}$ has $r_{+}$ hyperbolic nodes and $r_{-}$ elliptic nodes. In view of formulas (17), the nodes of $C_{3}$ yield the nodes
$(t_{1}(\lambda_{j}+O(t_{1})),\ t_{1}^{k}(\lambda_{j}^{k}+O(t_{1}))),\quad j=1,...,r\ ,$
of the curve ${\Phi}_{t}({\mathbb{P}}^{1})$, where $t_{1}=t-t^{*}$: for $t_{1}\lambda_{j}>0$ a hyperbolic node, for $t_{1}\lambda_{j}<0$ an elliptic one.
Thus, if $k$ is odd, then, for $t>t^{*}$ (respectively, $t<t^{*}$), the curve ${\Phi}_{t}({\mathbb{P}}^{1})$ gets $r_{+}$ (respectively, $r_{-}$) hyperbolic nodes in ${\operatorname{Tor}}(P_{\Delta})^{+}$ and $r_{-}$ (respectively, $r_{+}$) elliptic nodes outside ${\operatorname{Tor}}(P_{\Delta})^{+}$. None of these nodes contributes to the right-hand side of (8), and hence $W_{0}(\xi_{t})$ remains constant as $0<|t-t^{*}|<{\varepsilon}$.
If $k$ is even, then the curve ${\Phi}_{t}({\mathbb{P}}^{1})$ gets $r_{+}$ hyperbolic and $r_{-}$ elliptic nodes in ${\operatorname{Tor}}(P_{\Delta})^{+}$ if $t_{1}>0$, and gets $r_{-}$ hyperbolic and $r_{+}$ elliptic nodes in ${\operatorname{Tor}}(P_{\Delta})^{+}$ if $t_{1}<0$. Suppose that $r$ is even. Then $r_{+}\equiv r_{-}\mod 2$, and hence the first product in the right-hand side of formula (8) for $W_{0}(\xi_{t})$ remains constant. In turn, the orientation of the real local branch of ${\Phi}_{t}({\mathbb{P}}^{1})$ at $w$ remains the same (see Figure 2(a,b)), and hence the second product in the right-hand side of (8) remains constant as well. Suppose that $r$ is odd. Then $r_{+}\equiv r_{-}+1\mod 2$, and hence the first product in the right-hand side of formula (8) for $W_{0}(\xi_{t})$ changes sign as $t$ passes through $t^{*}$. However, the orientation of the real local branch of ${\Phi}_{t}({\mathbb{P}}^{1})$ at $w$ changes (see Figure 2(c,d)), which causes the change of sign in the second product in the right-hand side of (8).
[…]
Figure 2: Proof of Lemma 4.3, part (2), II
Thus, $W_{0}^{\kappa}(t)$ remains constant in the considered bifurcation.
(3) Suppose that $\xi$ is as in Lemma 4.2(2ii). Here we choose a path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ defined in a neighborhood of $t^{*}$ so that the point $w_{i}^{\sigma}(t)$ is fixed and the point $w_{j}(t^{*})$ is mobile (together with some other point of $\widehat{\boldsymbol{w}}(t)$). Since $C$ is immersed, we extract the required local constancy of $W_{0}^{\kappa}(t)$ from the transversality relation (12) and the smoothness of the two following strata in ${\mathcal{O}}_{C,w}/{\mathfrak{m}}_{w}^{s}$:
  • one stratum parameterizes deformations of a smooth branch intersecting ${\operatorname{Tor}}(\sigma)$ at $w$ with multiplicity $2(k_{i}^{\sigma}+k_{j}^{\sigma})$ into a smooth branch intersecting ${\operatorname{Tor}}(\sigma)$ at $w$ with multiplicity $2k_{i}^{\sigma}$ and in a nearby point with multiplicity $2k_{j}^{\sigma}$,
  • the other stratum parameterizes deformations of a couple of smooth branches intersecting ${\operatorname{Tor}}(\sigma)$ at $w$ with multiplicities $2k_{i}^{\sigma}$ and $2k_{j}^{\sigma}$, respectively, into a couple of smooth branches, one intersecting ${\operatorname{Tor}}(\sigma)$ at $w$ with multiplicity $2k_{i}^{\sigma}$ and the other intersecting ${\operatorname{Tor}}(\sigma)$ in a nearby point with multiplicity $2k_{j}^{\sigma}$.
Both smoothness statements are straightforward.
(4) Suppose that $\xi$ is as in Lemma 4.2(2iii) and ${\Phi}(p)\in{\operatorname{Tor}}_{\mathbb{R}}(P_{\Delta})^{\times}_{+}$. We closely follow the proof of [19, Theorem 5], providing here details for the reader’s convenience. Consider the path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ locally obtained from $\widehat{\boldsymbol{w}}(t^{*})$ by moving one point of $\widehat{\boldsymbol{w}}(t^{*})\cap C_{1}$ and one point of $\widehat{\boldsymbol{w}}(t^{*})\cap C_{2}$, while fixing the position of the other points of $\widehat{\boldsymbol{w}}(t^{*})$. Then, we obtain a smooth one-dimensional germ in ¯$\overline{\mathcal{M}}_{0,n}(\Delta)$, which locally induces a deformation of the hyperbolic node ${\Phi}(\widehat{C},p)$ equivalent in suitable local coordinates in a neighborhood of $z={\Phi}(p)$ to $uv=\lambda(t)$ with $\lambda$ a smooth function in a neighborhood of $t^{*}$ such that $\lambda(t^{*})=0$ and $\lambda^{\prime}(t^{*})>0$ (see Figure 3(a)). The smoothness and the local deformation claims are straightforward from the (standard) considerations in the proof of [12, Lemma 11(2)], which reduce both claims to the following $h^{0}$-vanishing (cf. [12, Formula (16) and computations in the first paragraph in page 251]):
$h^{0}(\widehat{C}_{s},{\mathcal{O}}_{\widehat{C}_{s}}({\boldsymbol{d}}_{s}))=0,\quad s=1,2\ ,$ (19)
where
$\deg{\boldsymbol{d}}_{s}=C_{s}^{2}-(C_{s}^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta}))[C_{s}]+2)-(c_{1}({\operatorname{Tor}}(P_{\Delta}))[C_{s}]-1)=-1>-2\ ,$
and hence (19) follows.
[…]
Figure 3: Proof of Lemma 4.3, part (4)
Now we analyze the change of the refined invariants. For, we pass to the consideration of complex oriented curves. Denote by $\widehat{C}_{1}^{\pm}$ and $\widehat{C}_{2}^{\pm}$ the connected components of $\widehat{C}_{1}\setminus{\mathbb{R}}\widehat{C}_{1}$ and $\widehat{C}_{2}\setminus{\mathbb{R}}\widehat{C}_{2}$, respectively. When $t$ varies around $t^{*}$, the curve $\widehat{C}_{1}\cup\widehat{C}_{2}$ turns into ${\mathbb{P}}^{1}$, and we encounter two types of deformations: either $\widehat{C}_{1}^{+},\widehat{C}_{1}^{-}$ glue up with $\widehat{C}_{2}^{+},\widehat{C}_{2}^{-}$, respectively, or $\widehat{C}_{1}^{+},\widehat{C}_{1}^{-}$ glue up with $\widehat{C}_{2}^{-},\widehat{C}_{2}^{+}$, respectively. The type of the deformation agrees with the local complex orientation of the central curve as shown in Figure 3(b). This means that if for $t<t^{*}$ one encounters a deformation of the first type, then for $t>t^{*}$ it is of the second type, and vice versa. Set (cf. (8))
[…]
Depending on the point of $C_{1}\cap C_{2}$ chosen for the smoothing, the invariant $(\kappa,{\varepsilon})$, where $\kappa$ is the quantum index and ${\varepsilon}$ is the second product in (8), changes as follows (see Figure 3(c)): either the real tangent vectors to $C_{1},C_{2}$ form a negative pair and
$\begin{cases}(\kappa_{1}+\kappa_{2},{\varepsilon}_{1}^{+}{\varepsilon}_{2}^{+}),&\\ (-\kappa_{1}-\kappa_{2},{\varepsilon}_{1}^{-}{\varepsilon}_{2}^{-}),&\end{cases}\ \Longrightarrow\ \begin{cases}(\kappa_{1}-\kappa_{2},{\varepsilon}_{1}^{+}{\varepsilon}_{2}^{-}),&\\ (-\kappa_{1}+\kappa_{2},{\varepsilon}_{1}^{-}{\varepsilon}_{2}^{+}),&\end{cases}$
or the real tangent vectors to $C_{1},C_{2}$ form a positive pair and
$\begin{cases}(\kappa_{1}-\kappa_{2},{\varepsilon}_{1}^{+}{\varepsilon}_{2}^{-}),&\\ (-\kappa_{1}+\kappa_{2},{\varepsilon}_{1}^{-}{\varepsilon}_{2}^{+}),&\end{cases}\ \Longrightarrow\ \begin{cases}(\kappa_{1}+\kappa_{2},{\varepsilon}_{1}^{+}{\varepsilon}_{2}^{+}),&\\ (-\kappa_{1}-\kappa_{2},{\varepsilon}_{1}^{-}{\varepsilon}_{2}^{-}).&\end{cases}$
For a simple geometric reason, the number of intersection points of the first kind equals the number of the intersection points of the second type. Hence, summing up over smoothings fo all real intersection points of $C_{1},C_{2}$, we derive the constancy of the numbers $W_{0}^{\kappa}(t)$ along the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$ ($0<\zeta\ll 1$) in case of ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ a birational immersion.
The case of multiple coverings can be treated in the same manner. We leave details to the reader.
(5) Suppose that $\xi$ is as in Lemma 4.2(2iii) and ${\Phi}(p)=w_{i}^{\sigma}(t^{*})=:w$. Observe that $({\Phi}\big{|}_{\widehat{C}_{s}})^{*}(w)=2l_{s}p$, $s=1,2$, where $l_{1}+l_{2}=k_{i}^{\sigma}$. Suppose that ${\Phi}\big{|}_{\widehat{C}_{1}},{\Phi}\big{|}_{\widehat{C}_{2}}$ both are immersions and that $l_{1}\leq l_{2}$.
Define the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$, $0<\zeta\ll 1$, by picking one mobile point in $C_{1}\cap\widehat{\boldsymbol{w}}(t^{*})\setminus\{w_{i}^{\sigma}(t^{*})\}$ and the other in $C_{2}\cap\widehat{\boldsymbol{w}}(t^{*})\setminus\{w_{i}^{\sigma}(t^{*})\}$. It defines a one-dimensional subvariety of ¯$\overline{\mathcal{M}}_{0,n}(\Delta)$, and let =${\mathcal{F}}=\{[{\Phi}_{t}:(\widehat{C}^{(t)},{\boldsymbol{p}}_{t})\to{\operatorname{Tor}}(P_{\Delta})]\}_{|t-t^{*}|<\zeta}$ be one of the irreducible germs of that variety at $\xi$, where $\widehat{C}^{(t)}\simeq{\mathbb{P}}^{1}$ as $t\neq 0$. The one-dimensional family of curves $C^{(t)}={\Phi}_{t}(\widehat{C}^{(t)})\in|{\mathcal{L}}_{P_{\Delta}}|$, $|t-t^{*}|<\zeta$, has a tangent cone at $C_{1}\cup C_{2}$ spanned by $C_{1}\cup C_{2}$ and some curve $C^{*}\in|{\mathcal{L}}_{P_{\Delta}}|\setminus\{C_{1}\cup C_{2}\}$.
[…]
$\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C_{1})}\delta(C_{1},z)+c_{1}({\mathcal{L}}_{P_{\Delta}})[C_{1}]-1=C_{1}^{2}+1\ .$
Further on, a monomial in an equation of $C^{*}$ below ${\Gamma}$ cannot be $x^{s}$, $0\leq s<2(l_{1}+l_{2})$, since $(C^{*}\cdot{\operatorname{Tor}}(\Sigma))_{w}\geq 2(l_{1}+l_{2})$, and it cannot be $x^{s}y$ with $20\leq s\leq 2l_{1}-$ 2. For the latter claim, we observe that each curve $C^{(t)}$, $t\neq 0$, has $12l_{1}-$ 1 nodes in a neighborhood of $w$ and satisfies $(C^{(t)}\cdot{\operatorname{Tor}}(\sigma))_{w}=2(l_{1}+l_{2})$. Thus, if the tangent line at $C^{(t)}$ to the considered family of curves is spanned by $C^{(t)}$ and $(C^{*})^{(t)}$, then $C^{(t)}$ and $(C^{*})^{(t)}$ intersect in a neighborhood of $w$ with multiplicity $\geq 2(2l_{1}-1)+2(l_{1}+l_{2})=4l_{1}+2l_{2}-2$. Hence, $C^{*}$ and $C_{1}\cup C_{2}$ intersect at $w$ with multiplicity $\geq 4l_{1}+2l_{2}-2$, which leaves the only possibility of the extra monomial $x^{2l_{1}-1}y$.
[…]
Figure 4: Proof of Lemma 4.3, part (5), I
If $l_{1}=l_{2}=l$, we can write down an equation of $C^{(t)}$ in a neighborhood of $w$ in the form
$y^{2}(1+O(t_{1}^{>0}))-(\eta_{1}+\eta_{2}+O(t_{1}^{>0}))x^{2l}y+x^{4l}(\eta_{1}\eta_{2}+O(t_{1}^{>0}))+\text{h.o.t.}+\sum_{r=0}^{2l-1}a_{r1}(t_{1})x^{r}y=0,$
where h.o.t. designates the terms above the Newton diagram, and
$t_{1}=t-t^{*},\quad a_{r1}(0)=0\quad\text{for all}\quad r=0,...,2l-1\ .$
Consider the tropical limit at $t_{1}\to 0$. It includes a subdivision of the triangle $T={\operatorname{conv}}\{(0,1),(4l,0),(02)\}$ and certain limit curves in the toric surfaces associated with the pieces of the subdivision. Since locally $C^{(t)}$ is an immersed cylinder, we obtain that the union of the limit curves is a curve of arithmetic genus zero, which finally allows the only following tropical limit: the triangle $T$ is the unique piece of the subdivision, and
$a_{r1}(t_{1})=t_{1}^{\lambda(2l-r)}(a_{r1}^{0}+O(t_{1}^{>0})),\quad r=0,...,2l-1\ ,$
where $a_{2l-1,1}^{0}\neq 0$ in view of the above conclusion on the Newton diagram of $C^{*}$, and in addition, the limit curve with Newton triangle $T$ given by
$y^{2}-(\eta_{1}+\eta_{2})x^{2l}y+\eta_{1}\eta_{2}x^{4l}+\sum_{r=0}^{2r-1}a_{r1}^{0}x^{r}y=0$ (20)
is rational. It also follows, that ${\mathcal{F}}$ is smooth, regularly parameterized by $t_{1}$ (i.e., $\lambda=1$), and its real part submersively projects onto the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$. By the patchworking theorem [21, Theorems 3.1 and 4.2], we uniquely restore ${\mathcal{F}}$ as long as we compute the coefficients $a_{r1}^{0}$, $20\leq r\leq 2l_{1}-$ 2 (here $a_{2l-1,1}^{0}$ is determined by $C^{*}$). The rationality of the curve (20) can be expressed as follows. Write equation (20) in the form $y^{2}-2P(x)y+\eta_{1}\eta_{2}x^{4l}=0$, then resolve with respect to $y$:
$y=P(x)\pm\sqrt{P(x)^{2}-\eta_{1}\eta_{2}x^{4l}}=P(x)\pm x^{2l}\sqrt{Q(1/x)^{2}-\eta_{1}\eta_{2}},\quad\deg Q=2l\ .$
The rationality means that the expression under the radical has $12l-$ 1 double roots (corresponding to the nodes of the curve). This means that $Q$ is a modified Chebyshev polynomial:
$Q(u)=\lambda_{1}{\operatorname{Cheb}}_{2l}(u+\lambda_{2})+\lambda_{3},\quad{\operatorname{Cheb}}_{2l}(u)=\cos(2l\cdot\arccos u)\ ,$
where $\lambda_{1},\lambda_{2},\lambda_{3}$ can be computed out of $\eta_{1},\eta_{2}$, and $a_{2l-1,1}^{0}$. As noticed in [22, Proof of Proposition 6.1], there are exactly two real solutions, one corresponding to a curve with the real part, shown in Figure 4(c) by lines inside the four Newton triangles designating four real quadrants, and this curve has one hyperbolic node and $22l-$ 2 non-real nodes, while the other curve, shown in Figure 4(d), has $12l-$ 1 elliptic nodes. Bringing the complex orientations to the play, we see that the former solution appears when the local real branches $({\mathbb{R}}C_{1},w)$ and ${\mathbb{R}}C_{2},w)$ are cooriented, while the latter solution corresponds to the case when these real branches have opposite orientations, see Figures 4(c,d). That is, on both sides of the path $\{{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$ we encounter four elements of $\overrightarrow{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))$ having quantum invariants
$\kappa_{1}+\kappa_{2},\ -\kappa_{1}-\kappa_{2},\ \kappa_{1}-\kappa_{2},\ \kappa_{2}-\kappa_{1}.$
It remains to notice that the Welschinger signs (8) of the elements with quantum indices $\kappa_{1}+\kappa_{2},-\kappa_{1}-\kappa_{2}$ are the same on both sides, since the modified limit curves shown in Figure 4(c) for $t>t^{*}$ and $t<t^{*}$ are obtained from each other by reflection with respect to the vertical axis. The same holds for the elements with quantum indices $\kappa_{1}-\kappa_{2},\kappa_{2}-\kappa_{1}$ (see Figure 4(d)). This completes the proof of the constancy of the numbers $W_{0}^{\kappa}(t)$ in the considered case $l_{1}=l_{2}$.
In the case $l_{1}<l_{2}$, an equation of $C^{(t)}$ in a neighborhood of $w$ takes the form
[…]
where $t_{1}=t-t^{*}$, $a_{r1}(0)=0$ for all $r=0,...,2l-1$, and h.o.t. stands for the terms above the Newton diagram. In this situation, due to the condition of arithmetic genus zero, the tropical limit is defined uniquely: the area under the Newton diagram is divided into two triangles
${\operatorname{conv}}\{(0,1),(2l_{1},1),(2(l_{1}+l_{2}),0)\}\quad\text{and}\quad{\operatorname{conv}}\{(0,1),(2l_{1},0),(0,2)\}$
(see Figure 5(a)), while the monomials on the segment $[(0,1),(2l_{1},1)]$ sum up to $\eta_{1}(x+\lambda t_{1})^{2l_{1}}y$, where $\lambda\neq 0$ is uniquely determined by the coefficient $a_{2l_{1}-1,1}^{0}$ coming from $C^{*}$. The two limit curves have branches intersecting each other with multiplicity $2l_{1}$ along the toric divisor ${\operatorname{Tor}}([(0,1),(2l_{1},0)])$ (see Figure 5(b)). The genuine geometry of $C^{(t)}$ can be recovered when we deform that intersection point into $12l_{1}-$ 1 nodes. The modification describes such a deformation as a replacement of the intersection point by one of the $2l_{1}$ modified limit curves, among which exactly two are real, and their real parts are shown in Figure 5(c,d) (cf. Figures 4(c,d)): one of them has a hyperbolic node and $22l_{1}-$ 2 non-real nodes, and the other $12l_{1}-$ 1 elliptic nodes. As in the preceding paragraph, the former modified limit curve fits the case when the local real branches $({\mathbb{R}}C_{1},w)$, $({\mathbb{R}}C_{2},w)$ are cooriented, which means that on each side of the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$ we have two elements of $\overrightarrow{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))$ with quantum indices $\kappa_{1}+\kappa_{2},-\kappa_{1}-\kappa_{2}$, while all four elements have the same Welschinger sign (8) since the constructions for $t>t^{*}$ and $t<t^{*}$ are symmetric with respect to the vertical axis. Hence, the constancy of the numbers $W_{0}^{\kappa}(t)$. The same holds in the case of the opposite orientation of the branches $({\mathbb{R}}C_{1},w)$, $({\mathbb{R}}C_{2},w)$ with the use of the second modified limit curve. We also point out that the final deformation including the modification still depends on $t_{1}=t-t^{*}$.
Refer to caption(a)(b)(c)(d)$2l_{1}$$2l_{1}+2l_{2}$11112222
Figure 5: Proof of Lemma 4.3, part (5), II
The remaining case is as follows: ${\Phi}:\widehat{C}_{1}\to{\operatorname{Tor}}(P_{\Delta})$ is an $2l_{1}$-multiple covering of a line through $w$ and ${\Phi}:\widehat{C}_{2}\to{\operatorname{Tor}}(\Sigma)$ is an immersion onto a rational curve having a smooth branch at $w$ intersecting ${\operatorname{Tor}}(\sigma)$ with multiplicity $2l_{2}$. Here we remove the restriction $l_{1}\leq l_{2}$. We extend the range of the coordinate system $x,y$ to a neighborhood of the line $C_{1}={\Phi}(\widehat{C}_{1})$, assuming that the axis $\{x=0\}$ is the whole line $C_{1}$. Then the Newton diagram of ${\Phi}_{*}(\widehat{C})$ in this neighborhood of $C_{1}$ is as shown in Figure 6(a) by fat lines, where $s=C_{1}C_{2}$. Following the recipe used in the preceding cases, we obtain that the tropical limit of the family $\{C^{(t)}\}_{|t-t^{*}|<\zeta}$ as $t_{1}:=t-t^{*}\to 0$ includes the triangle ${\operatorname{conv}}\{(0,1),(2l_{1},1),(2(l_{1}+l_{2}),0)\}$ and the rectangle ${\operatorname{conv}}\{(0,1),(0,s+1),(2l_{1},1),(2l_{1},s+1)\}$ (see Figure 6(b)), and, in addition, the part of the equation of $C^{(t)}$ restricted to the union of these polygons is as follows:
$y\cdot((x-\lambda_{1}t_{1})^{2l_{1}}-a(x-\lambda_{2}t_{1})^{2l_{1}}y)\prod_{j=1}^{s-1}(y-y_{j})-\eta x^{2(l_{1}+l_{2})}=0\ ,$
where $y_{1},...,y_{s-1}$ are the coordinates of the points $C_{1}\cap C_{2}\setminus\{w\}$, $\lambda_{1}\neq 0$ is determined by $C^{*}$, $\lambda_{2}t_{1}\neq 0$ is the current coordinate of the mobile point from $C_{1}\cap\widehat{\boldsymbol{w}}(t^{*})\setminus\{w\}$, and $a,\eta>0$ are determined by $C_{2}$. Again the two limit curves have intersection of multiplicity $2l_{1}$ in a point on the toric divisor ${\operatorname{Tor}}([(0,1),(2l_{1},1)])$. The genuine geometry of $C^{(t)}$ in a neighborhood of $w$ is obtained by deforming this tangency point by means of the two real modified limit curves, which we used in the preceding paragraph. Then, the argument from the preceding paragraph in the same manner completes the proof of the constancy of the numbers $W_{0}^{\kappa}(t)$ in the considered wall-crossing bifurcation.
(5b) The fact that the (global) family ${\mathcal{F}}$ is smooth, realizes the local wall crossing described in part (5a), and transversally crosses the wall amounts by the standard argument used in part (4), to the condition (19). The latter condition can be verified exactly as in part (4).
(6) Suppose that $\xi$ is as in Lemma 4.2(2iv). Without loss of generality we can suppose that
$P={\operatorname{conv}}\{(0,2m),\ (2p,0),\ (2q,0)\},\quad 0\leq p<q,\ m<q\ ,$
and that ${\Phi}$ is ramified at the points of $\widehat{\boldsymbol{w}}(t^{*})$ on the toric divisors associated with the segments $[(0,2m),(2p,0)]$ and $[(0,2m),(2q,0)]$, while the two points $w_{1}^{\sigma}(t),w_{2}^{\sigma}(t)\in{\boldsymbol{w}}(t)$ on the toric divisor ${\operatorname{Tor}}(\sigma)$, where $\sigma=[(2p,0),(2q,0)]$, merge to one point as $t\to t^{*}$. Then $C={\Phi}({\mathbb{P}}^{1})$ admits a parametrization
$x=a\theta^{m},\quad y=b\theta^{p}(\theta-1)^{q-p},\quad\theta\in{\mathbb{C}}\ ,$
and correspondingly ${\Phi}$ is given by
$x=a\theta^{2m},\quad y=b\theta^{2p}(\theta^{2}-1)^{q-p},\quad\theta\in{\mathbb{C}}\ .$
Assuming that the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<{\varepsilon}}$, is defined by fixing the point $w_{1}^{\sigma}$, associated with $\theta=1$, and the point of $\widehat{\boldsymbol{w}}(t^{*})$ on ${\operatorname{Tor}}([(0,2m),(2q,0)])$, we obtain a one-parameter deformation of ${\Phi}$
$x=a\theta^{2m},\quad y=b\theta^{2p}(\theta-1)^{q-p}(\theta+1+\lambda)^{q-p},\quad|\lambda|\ll 1\ .$ (21)
This deformation is regularly parameterized by the difference of the $x$-coordinates of $w_{2}^{\sigma}(t)$ and $w_{1}^{\sigma}(t)$ equal to $2ma\lambda+O(\lambda^{2})$. It follows from formulas (21) that the corresponding element $\xi(t)\in\overline{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))$ has the same orientations at a point on the toric divisor ${\operatorname{Tor}}([(0,2m),(2p,0)])$ for $\lambda<0$ and for $\lambda>0$, and so does for the point on the toric divisor ${\operatorname{Tor}}([(0,2m),(2q,0)])$, while the local branches at $w_{1}^{\sigma}(t)$, $w_{2}^{\sigma}(t)$ have opposite orientations with respect to the orientation of ${\operatorname{Tor}}_{\mathbb{R}}(\sigma)$. Furthermore, $\xi(t)$ has exactly two elliptic nodes in a neighborhood of each elliptic node of $C$. Thus, the constancy of $W_{0}^{\kappa}(t)$ follows. $\Box$
[…]
Figure 6: Proof of Lemma 4.3, part (5), III
5 Invariance in genus one: proof of Theorem 3.7
Take two sequences $\widehat{\boldsymbol{w}}(0)=(\widehat{\boldsymbol{w}}_{\partial}(0),w_{0}(0))$ and $\widehat{\boldsymbol{w}}(1)=(\widehat{\boldsymbol{w}}_{\partial}(1),w_{0}(1))$, satisfying the conditions of Theorem 3.7. We may assume that $\widehat{\boldsymbol{w}}_{\partial}(0),\widehat{\boldsymbol{w}}_{\partial}(1)\in\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ for some $\tau>0$. Then, we join these sequences by a generic path $\{\widehat{\boldsymbol{w}}_{\partial}(t)\}_{0\leq t\leq 1}$, in $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ and the points $w_{0}(0)$, $w_{0}(1)$ by a generic path $\{w_{0}(t)\}_{0\leq t\leq 1}$ in ${\mathcal{Q}}$, and verify the constancy of $W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ in all possible wall-crossing events, for all $\kappa$.
5.1 Invariance inside chambers
Similarly to the proof of Theorem 3.2, we start with the verification of the constancy of $W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ along intervals $t^{\prime}<t<t^{\prime\prime}$ such that $\{\widehat{\boldsymbol{w}}(t)\}_{t^{\prime}<t<t^{\prime\prime}}$ lies in the same chamber of $\operatorname{Men}_{2}(\Delta)$. To this end, it is enough to show that the projection ${\mathcal{M}}_{1,n}^{\;{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}\to(t^{\prime},t^{\prime\prime})$ has no critical points (cf. the proof of Theorem 3.2). This requirement amounts to the following statement (cf., Lemma 4.1 and [23, Lemma 2.3]).
Lemma 5.1
Let $\xi=[{\Phi}:({\mathbb{E}},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{1,n}^{\;{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})$, where ${\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)$. Then
$H^{0}\left({\mathbb{E}},{\mathcal{N}}_{\Phi}\left(-p_{0}-\sum_{\sigma\neq\sigma_{0},i\neq i_{0}}2k^{\sigma}_{i}p^{\sigma}_{i}-(2k^{\sigma_{0}}_{i_{0}}-1)p^{\sigma_{0}}_{i_{0}}\right)\right)=0.$ (22)
Proof. We proceed along the lines of the proof of [23, Lemma 2.3]. Assume that $H^{0}\neq 0$ in (22). Then, there exists a real curve $C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|\setminus\{C\}$, where $C={\Phi}_{*}({\mathbb{E}})$, which intersects $C$
  • at each point $q\in{\operatorname{Sing}}\hskip 1.70709pt(C)$ with multiplicity $\geq 2\delta(C,q)$,
  • at each point $w^{\sigma}_{i}$, $(\sigma,i)\neq(\sigma_{0},i_{0})$ with multiplicity $\geq 2k^{\sigma}_{i}$,
  • at $w^{\sigma_{0}}_{i_{0}}$ with multiplicity $\geq 2k^{\sigma_{0}}_{i_{0}}-1$,
  • and at $w_{0}$.
Since $C$ has in $\overline{\mathcal{Q}}$ a null-homologous immersed circle $S$, there must be an additional intersection point $w^{\prime}\in C^{\prime}\cap S$, not mentioned in the above list. However, then
$C^{\prime}C\geq 2\sum_{q\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,q)+\sum_{\sigma\in P^{1}_{\Delta}}\sum_{i=1}^{n^{\sigma}}2k^{\sigma}_{i}-1+2$
$=(C^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta})[C]))+c_{1}({\operatorname{Tor}}(P_{\Delta})[C]-1+2=C^{2}+1,$
which is a contradiction. $\Box$
5.2 Walls in the space of constraints
The set of sequences $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)\times{\mathcal{Q}}$ satisfying
¯$\overline{\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})={\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})\quad\text{and}\quad\widehat{\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta),$
is a dense semialgebraic subset of full dimension $\dim\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)\times{\mathcal{Q}}=n+1$. The complement is the union of finitely many semialgebraic strata of codimension $\geq 1$. Since the path $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{0\leq t\leq 1}$ is generic, it avoids strata of $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)\times{\mathcal{Q}}$ of codimension $\geq 2$ and intersects strata of codimension one only in their generic points.
Now, we study the strata $S$ of the dimension $n$ in $(\operatorname{Men}_{\mathbb{R}}^{\tau}\times{\mathcal{Q}})\setminus\operatorname{Men}_{1}(\Delta)$ in terms of the geometry of elements $\xi\in\overline{\mathcal{M}}_{1,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}}(t))\setminus{\mathcal{M}}_{1,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}}(t))$, $0\leq t\leq 1$.
Lemma 5.2
(1) The following elements $\xi=[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]$ cannot occur in ¯$\overline{\mathcal{M}}_{1,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}}(t))\setminus{\mathcal{M}}_{1,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}}(t))$, $0\leq t\leq 1$, as $\widehat{\boldsymbol{w}}\in S$:
  1. (1i)
    $\widehat{C}$ is a reducible curve of arithmetic genus $1$ with a component mapped onto a toric divisor;
  2. (1ii)
    $\widehat{C}$ is a connected curve of arithmetic genus $1$ either with at least three irreducible components, or with two rational irreducible components and no other components.
(2) If $\xi=[{\Phi}:(\widehat{C},\widehat{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in\overline{\mathcal{M}}_{1,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$, where $\widehat{\boldsymbol{w}}$ is a generic element in an $n$-dimensional stratum in $(\operatorname{Men}_{\mathbb{R}}^{\tau}\times{\mathcal{Q}})\setminus\operatorname{Men}_{1}(\Delta)$, then $\xi$ is of one of the following types:
(2a) either $w_{0}(t^{*})$ is a singular point of the curve $C={\Phi}_{*}(\widehat{C})$, and the following holds:
  1. (2i)
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, the curve $\widehat{C}$ is smooth elliptic, ${\Phi}$ is an immersion onto a curve $C$ that is smooth along the toric divisors, while $w_{0}(t^{*})$ is a center of at least one real local branch;
(2b) or $w_{0}(t^{*})$ is a smooth point of the curve $C={\Phi}(\widehat{C})$, and one of the following holds:
  1. (2ii)
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, the curve $\widehat{C}$ is smooth elliptic, ${\Phi}$ is birational onto its image, but not an immersion; furthermore, $C={\Phi}(\widehat{C})$ is smooth at $w_{0}(t^{*})$ and at each point $w_{i}^{\sigma}(t^{*})$ with $k_{i}^{\sigma}\geq 2$, and is unibranch at each point $w_{i}^{\sigma}(t^{*})$ with $k_{i}^{\sigma}=1$;
  2. (2iii)
    $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})$ for some $\sigma\in P^{1}_{\Delta}$ and $i\neq j$, the curve $\widehat{C}$ is smooth elliptic, and the map ${\Phi}$ is an immersion such that the point $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})$ is a center of one or two smooth branches;
  3. (2iv)
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $\widehat{C}_{1}\simeq{\mathbb{P}}^{1}$ and $\widehat{C}_{2}$ is a smooth elliptic curve, the intersection $\widehat{C}_{1}\cap\widehat{C}_{2}$ consisting of one point $p$; the map ${\Phi}:\widehat{C}_{1}\to{\operatorname{Tor}}(P_{\Delta})$, $s=1,2$, is either an immersion, smooth along ${\operatorname{Tor}}(\partial P_{\Delta})$, or a multiple covering of a line intersecting only two toric divisors, while these divisors correspond to opposite parallel sides of $P_{\Delta}$ and the intersection points with these divisors are ramification points of the covering; the map ${\Phi}:\widehat{C}_{2}\to{\operatorname{Tor}}(P_{\Delta})$ is an immersion, smooth along ${\operatorname{Tor}}(\partial P_{\Delta})$; the point $p$ either is mapped to ${\operatorname{Tor}}(P_{\Delta})^{\times}$, and then the curves $C_{1}={\Phi}(\widehat{C}_{1})$, $C_{2}={\Phi}(\widehat{C}_{2})$ intersect only in ${\operatorname{Tor}}(P_{\Delta})^{\times}$ and each of their intersection point is an ordinary node, or $p$ is mapped to some point $w_{i}^{\sigma}(t^{*})$, in which case the curves $C_{i}={\Phi}(\widehat{C}_{i})$, $i=1,2$, intersect ${\operatorname{Tor}}(\sigma)$ with even multiplicities, and they do not have other common point in $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$;
  4. (2v)
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, $\widehat{C}$ is an irreducible rational curve with one node $p$, the map ${\Phi}$ is an immersion that sends $p$ to some point $w_{i}^{\sigma}(t^{*})$, center of two smooth branches intersecting ${\operatorname{Tor}}(\sigma)$ with odd multiplicities $12l_{1}+1,2l_{2}+$ 1, where $l_{1}+l_{2}+1=k_{i}^{\sigma}$, and intersecting each other with multiplicity $\min\{2l_{1}+1,2l_{2}+1\}$, furthermore, each of the other points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ is a center of one smooth branch;
  5. (2vi)
    $\widehat{C}$ is a smooth elliptic curve, $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n=4$ points, the curve $C={\Phi}({\mathbb{P}}^{1})$ is rational and smooth at $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, and ${\Phi}:\widehat{C}\to C$ is a double covering ramified at $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$.
Proof. (1i) In such a case, we must have $\widehat{C}=\widehat{C}^{\prime}\cup\widehat{C}^{\prime\prime}$, where $\widehat{C}^{\prime}$ is a connected curve of arithmetic genus one suhc that $\widehat{C}^{\prime}$ is mapped onto the union of all toric divisors, and $\widehat{C}^{\prime\prime}$ is a non-empty (due to $w_{0}\in{\operatorname{Tor}}(P_{\Delta})^{\times}$) union of connected curves of arithmetic genus zero, each one joined with $\widehat{C}^{\prime}$ in one point (cf. the proof of Lemma 4.2(1i)). Let $\widehat{C}^{\prime\prime}_{0}$ be a connected component of $\widehat{C}^{\prime\prime}$ such that $w_{0}(t^{*})\in{\Phi}(\widehat{C}^{\prime\prime}_{0})$. Note that all local branches of ${\Phi}:\widehat{C}^{\prime\prime}_{0}\to{\operatorname{Tor}}(P_{\Delta})$ centered on toric divisors, except for ${\Phi}:(\widehat{C}^{\prime\prime}_{0},p)\to{\operatorname{Tor}}(P_{\Delta})$, where $p=\widehat{C}^{\prime}\cap\widehat{C}^{\prime\prime}_{0}$, are, in fact, centered at some points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$. Observe that either $C^{\prime\prime}_{0}$ intersects at least three toric divisors of ${\operatorname{Tor}}(P_{\Delta})$, or it intersects two toric divisors corresponding to parallel sides of $P_{\Delta}$. However, in the both cases we reach a contradiction with (AQC).
(1ii) In the case of at least $3$ irreducible components, we follow the lines of the proof of Lemma 4.2(1ii). If $\widehat{C}$ contains an elliptic component and at least $2$ rational components, then we similarly obtain at least three independent Menelaus type conditions for $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, which bounds from above the dimension of the considered stratum by $(n-3)+2=n-1<n$, a contradiction. If all irreducible components of $\widehat{C}$ are rational, then we encounter at least two independent Menelaus conditions on $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ as well as a condition on the position of the point $w_{0}(t^{*})$ due to the finiteness statement of Lemma 2.4, which altogether bounds from above the dimension of the considered stratum by $(n-2)+1=n-1<n$.
(2a) Let $w_{0}(t^{*})$ be a singular point of $C$. Then, for the dimension reason, $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ must be a generic element of $\operatorname{Men}_{\mathbb{R}}^{\tau}$. This implies, in particular, that $\widehat{C}$ is irreducible and $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points. Furthermore, $\widehat{C}$ cannot be a rational curve with a node. Thus, $\widehat{C}$ is a smooth elliptic curve. Then (see Lemma 2.4) we derive that ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ is an immersion and $C$ is smooth along the toric divisors. Since $w_{0}(t^{*})$ turns into a real smooth point in the deformation along the path $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{0\leq t\leq 1}$, it must be a center of at least one real local branch.
(2b) From now on we can assume that $w_{0}(t^{*})$ is a smooth point of $C={\Phi}_{*}(\widehat{C})$.
Suppose that $\widehat{C}$ is a smooth elliptic curve and $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points. Then, we obtain the claim (2ii) due to Lemma 2.4 and the claim (1ii) above.
Suppose that some of the points of the sequence $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ coincide. For the dimension reason, we immediately get that $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points, and all these points must be in general position subject to the unique Menelaus relation (1). Hence, $\widehat{C}$ must be irreducible. Moreover, $\widehat{C}$ cannot be rational, since otherwise, by Lemma 2.4, one would get the total dimension of the considered stratum $\leq(n-2)+1=n-1<n$, a contradiction. Thus, $\widehat{C}$ is a smooth elliptic curve. We have $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})=w$ for some $\sigma\subset\partial P$ and $i\neq j$. If $C$ is unibranch at each point of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, then, by Lemma 2.4, the curve $C$ is immersed and smooth along the toric divisors. Otherwise, the curve $C$ has two local branches at $w$. Moreover, it must be unibranch at each point $w_{i^{\prime}}^{\sigma^{\prime}}(t^{*})\neq w$ and, in addition, smooth if $k_{i^{\prime}}^{\sigma^{\prime}}\geq 2$ due to Lemma 2.4 and the claim (1ii). Let us show that $C$ is immersed. Fixing the position of $w_{0}(t^{*})$ and the position of one more point $w^{\prime}\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$, we obtain a family of dimension $\geq n-2\geq 2$; hence, [13, Inequality (5) in Lemma 2.1] applies:
$c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})\geq(c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})-n+2)+1+\sum_{B}({\operatorname{ord}}B-1)+(n-3)$
$=c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})+\sum({\operatorname{ord}}B-1)\ ,$
where $B$ runs over all singular local branches of $C$ in ${\operatorname{Tor}}(P_{\Delta})^{\times}\cup\{w,w^{\prime}\}$. That is, $C$ is immersed, and we finally fit the requirements of claim (2iii).
Thus, we are left with the case of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consisting of $n$ distinct points and $\widehat{C}$ either consisting of two components, or being a rational curve with a node. Observe that the case of two rational components is not possible, since by Lemma 2.4, the dimension of such a stratum would not exceed $(n-2)+1=n-1<n$.
Suppose that $\widehat{C}$ is as in item (2iv). For the dimension reason, the sequence $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ is in general position subject to exactly two Menelaus type relations, while $w_{0}(t^{*})$ is in general position in ${\mathcal{Q}}$. Thus, by Lemma 2.4, we obtain the immersion and smoothness statements as required. The rest of the argument literally coincides with the corresponding part of the proof of Lemma 4.2(2iii).
Suppose that $\widehat{C}$ is a rational curve with a node $p$. Then, this node cannot be mapped to ${\operatorname{Tor}}(P_{\Delta})^{\times}$, since otherwise one would encounter a real rational curve with two one-dimensional branches, one in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ and the other in $\overline{\mathcal{Q}}$. Hence the node is mapped to $w\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, and for the above reason, the intersection multiplicities of the two branches of $C={\Phi}(\widehat{C})$ at $w$ are odd. Observe that ${\Phi}^{\prime}:{\mathbb{P}}^{1}\to\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ is not a multiple covering of the image. Indeed, in such a case, one would have that $P$ is a triangle, $l_{1}=l_{2}=l$, and $C={\Phi}(\widehat{C})$ is a rational curve intersecting at least two of the toric divisors with odd multiplicity, the double covering being ramified at the two points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$. With an automorphism of ${\mathbb{Z}}^{2}$, we can turn the Newton triangle of $C$ (which is $\frac{1}{2}P_{\Delta}$) into the triangle
${\operatorname{conv}}\{(0,s),(r,0),(r+2l+1,0)\},\quad\gcd(s,r)\equiv 1\mod 2\ ,$
with $w\in{\operatorname{Tor}}([(r,0),(r+2l+1,0)])$ and $C$ given by a parametrization (in some affine coordinates $x,y$)
$x=a\theta^{s},\quad y=b\theta^{r}(\theta-1)^{2l+1},\quad a,b\in{\mathbb{R}}^{\times}\ .$
Since $x(1)>0$, we get $a>0$, and hence $x>0$ as $0<\theta\ll 1$. Then $y>0$ for $0<\theta\ll 1$, and hence $b<0$. Thus, if $r$ and $s$ are odd, we obtain $x>0,y<0$ as $\theta\to+\infty$, and $x<0,y<0$ as $\theta\to-\infty$. If $r$ is even and $s$ is odd, we get $x>0,y<0$ as $\theta\to+\infty$, and $x<0,y>0$ as $\theta\to-\infty$. In both cases, it follows that the intersection point with the toric divisor ${\operatorname{Tor}}([(0,s),(r+2l+1,0)])$ is out of $\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ against the initial assumption. If $r$ is odd and $s$ is even, we get that the segment of ${\mathbb{R}}C$ in the quadrant $\overline{\{x>0,y<0\}}$ touches all three toric divisors, which contradicts the condition (AQC). Thus, a multiple covering is excluded. The rest of claims can be derived in the same way as the statement of Lemma 4.2(2ii) with one exception: we shall show that the local branches of $C$ at $w$ intersect each other with multiplicity $\min\{2l_{1}+1,2l_{2}+1\}$. For $l_{1}\neq l_{2}$, this is immediate. Suppose that $l_{1}=l_{2}=l$. If $n=3$, then $P$ is a triangle and, in the above setting, we get the parametrization of $C$ in the form
$x=a\theta^{2s},\quad y=b\theta^{2r}(\theta-1)^{2l+1}(\theta-\theta_{0})^{2l+1},\quad\theta_{0}\neq 1,\quad a,b\in{\mathbb{R}}^{\times}\ .$
Since $x(1)=x(\theta_{0})$, we obtain the only real solution $\theta_{0}=-1$, and hence $y=b\theta^{2r}(\theta^{2}-1)^{2l-1}$ is a function of $\theta^{2}$ as well as $x$ is, which means that this is a double covering, forbidden above.
Suppose now that $\widehat{C}$ is a smooth elliptic curve and ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ is a multiple covering of its image. Since the preimages of at least $n-2$ points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ in $\widehat{C}$ are irreducible, the map ${\Phi}:\widehat{C}\to C\subset{\operatorname{Tor}}(P_{\Delta})$ is ramified, and hence $C$ cannot be elliptic by the Riemann-Hurwitz formula. Thus, $C$ is rational, which for dimension reason implies that $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ points in general position subject to one Menelaus condition. In particular, $C$ is smooth at each point of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, and ${\Phi}$ has ramification index $s$ at all points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$. By the Riemann-Hurwitz formula,
$0\leq 2s-(s-1)n.$
[…]
5.3 Wall-crossing
We complete the proof of Theorem 3.7 with the following lemma.
Lemma 5.3
Let $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{0\leq t\leq 1}$ be a generic path in $\operatorname{Men}_{\mathbb{R}}^{\tau}\times{\mathcal{Q}}$, and let $t^{*}\in(0,1)$ be such that ¯$\overline{\mathcal{M}}_{1,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ contains an element $\xi$ as described in one of the items of Lemma 5.2(2). Then, for each $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq{\mathcal{A}}(\Delta)$, the numbers $W_{1}^{\kappa}(t):=W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ do not change as $t$ varies in a neighborhood of $t^{*}$.
Proof. We always can assume that, in a neighborhood of $t^{*}$, the path $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{0\leq t\leq 1}$ is defined by fixing the position of $w_{0}(t^{*})$ and some $n-2$ points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, while the other two points remain mobile. Except for the case of Lemma 5.2(2iv) describing reducible degenerations, we work with families of curves which are trivially covered by families of complex oriented curves so that the quantum index persists along each component of the family of oriented curves.
(1) Suppose that $\xi$ is as in Lemma 5.2(2i). Consider the lifts of $w_{0}(t^{*})$ upon ${\mathbb{E}}$ that correspond to real branches of $C$ centered at $w_{0}(t^{*})$. Since $C$ is immersed, we only have to show that each of the lifts induces a germ of a smooth one-dimensional subfamily in ${\mathcal{M}}_{1,n}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$. This follows from the smoothness statements in [3, Proposition 4.17(2)] and in [24, Lemma 3(1)], [13, Lemma 2.4(2)].
(2) Suppose that $\xi$ is as in Lemma 5.2(2ii). The argument used in step (1) of the proof of Lemma 4.3 applies, provided we establish a relevant transversality statement. Here, the transversality condition is similar to (12):
$h^{0}({\mathbb{E}},{\mathcal{O}}_{\mathbb{E}}({\boldsymbol{d}}-p_{0}))=0,\quad{\boldsymbol{d}}\in{\operatorname{Div}}({\mathbb{E}})\ ,$ (23)
where
$\deg{\boldsymbol{d}}=C^{2}-\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}2\delta(C,z)-\sum_{w\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})}\dim{\mathcal{O}}_{C,w}/I_{w}$
$=C^{2}-(C^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}}))-(c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})-1)=1>0\ .$
Observe that by the Riemann-Roch formula, $h^{0}({\mathbb{E}},{\mathcal{O}}_{\mathbb{E}}({\boldsymbol{d}}))=1$, and hence (23) holds due to the general position of $w_{0}$ on $C$.
(3) The treatment of the case where $\xi$ is as in Lemma 5.2(2iii) literally coincides with that in step (2) of the proof of Lemma 4.3, when we replace the transversality relation (12) with (23).
(4) Suppose that $\xi$ is as in Lemma 5.2(2iv). Due to the condition (AQC), the real part of the rational curve $C_{1}={\Phi}(\widehat{C}_{1})$ must lie in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$, and the real part of the elliptic curve $C_{2}={\Phi}(\widehat{C}_{2})$ has two one-dimensional components, one in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ and the other in $\overline{\mathcal{Q}}$. Then, the local argument proving the constancy of the numbers $W_{1}^{\kappa}(t)$ literally coincides with that in step (5a) of the proof of Lemma 4.3.
The global argument is similar to that in step (5b) of the proof of Lemma 4.3. Here, we obtain the conditions
[…]
and
[…]
The former one follows by the Riemann-Roch formula, and the latter one as in part (2) above.
(5) Suppose that $\xi$ is as in Lemma 5.2(2v). Since $\widehat{C}$ is rational, both local branches $B_{1},B_{2}$ of $C={\Phi}(\widehat{C})$ at $w=w_{i}^{\sigma}(t^{*})$ are real and each one intersects ${\operatorname{Tor}}(\sigma)$ with odd multiplicities $12l_{1}+1,2l_{2}+$ 1, respectively, where $2l_{1}+2l_{2}+2=2k_{i}^{\sigma}$, $l_{1}\leq l_{2}$. Choose two mobile points $w_{i^{\prime}}^{\sigma^{\prime}},w_{i^{\prime\prime}}^{\sigma^{\prime\prime}}\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$. The one-dimensional stratum {¯$\{\overline{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,{\boldsymbol{w}}(t))\}_{|t-t^{*}|<\zeta}$, $0<\zeta\ll 1$, projects to the germ at $C$ of a one-dimensional variety in $|{\mathcal{L}}_{P_{\Delta}}|$, and denote by ${\mathcal{V}}$ one of its real irreducible components. Let the tangent line to ${\mathcal{V}}$ at $C$ be spanned by $C$ and $C^{*}\in|{\mathcal{L}}_{P_{\Delta}}|\setminus\{C\}$. Fix equavariant coordinates $x,y$ in a neighborhood of $w$ so that $w=(0,0)$ and ${\operatorname{Tor}}(\sigma)=\{y=0\}$. Then, $C$ in a neighborhood of $w$ is given by an equation
$(y+\eta_{1}x^{2l_{1}+1})(y+\eta_{2}x^{2l_{2}+1})+\text{h.o.t.}=0,\quad\eta_{1},\eta_{2}\in{\mathbb{R}}^{\times}\ ,$
with the Newton diagram (see Figure 1(a,b))
${\Gamma}=[(0,2),(2l_{1}+1,1)]\cup[(2l+1,1),(2l_{1}+2l+2+2,0)]$
(cf. Figures 5(a,b)). Exactly as in step (5a) of the proof of Lemma 4.3, an equation of $C^{*}$ has the monomial $x^{2l_{1}}y$ with a non-zero coefficient and no other monomials below the Newton diagram ${\Gamma}$.
If $l_{1}=l_{2}=l$, then as in step (5a) of the proof of Lemma 4.3, we obtain that the equation of the germ of $C^{(t)}\in{\mathcal{V}}$ at $w=(0,0)$ is as follows:
$y^{2}+P(x)y+\eta_{1}\eta_{2}x^{4l+2}+\text{h.o.t.}=0,\quad P(x)=(\eta_{1}+\eta_{2})x^{2l+1}+\sum_{j=2l}^{0}a_{j}t_{1}^{2l+1-j}x^{j}\ ,$ (24)
$\text{where}\quad t_{1}=t-t^{*},\quad\text{and}\quad x^{2l+1}P(1/x)=\lambda_{1}{\operatorname{Cheb}}_{2l+1}(t_{1}x+\lambda_{2})+\lambda_{3}$
with real numbers $\lambda_{1},\lambda_{2},\lambda_{3}$ uniquely determined by $\eta_{1},\eta_{2}$ and the coefficient of $x^{2l}y$ in the equation of $C^{*}$, and h.o.t. includes higher powers of $t_{1}$ in the present monomials and the sum of monomials above ${\Gamma}$. If $\eta_{1}\eta_{2}>0$, the latter formula yields $2l$ elliptic nodes in a neighborhood of $w$. Observe that the change of sign of $t_{1}$ yields the symmetry with respect to the origin for the real part of $C^{(t)}$ (up to h.o.t.), and hence (see Figure 1(c)) the quantum index and the Welschinger sign (9) persists along the path $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{|t-t^{*}|<\zeta}$. If $\eta_{1}\eta_{2}<0$, then equation (24) implies
[…]
which means that $C^{(t)}$ has only non-real nodes in a neighborhood of $w$. Again (see Figure 1(d)), the quantum index and the Welschinger sign remain unchanged.
[…]
Figure 7: Proof of Lemma 5.3, part (5)
Basically, the same local bifurcation occurs when If $l_{1}<l_{2}$ (cf. step (5a) of the proof of Lemma 4.3). The family ${\mathcal{V}}$ appears to be smooth, regularly parameterized by $t_{1}=t-t^{*}$ in the form
[…]
where $\eta_{0}$ is determined by $C^{*}$, and h.o.t. includes higher powers of $t_{1}$ in the present monomials and the sum of monomials above ${\Gamma}$. We skip the rest of details.
Remark 5.4
In equations (24) and (25), the sign of the coefficient of $x^{2l_{1}+2l_{2}+2}$ remains constant, and the sign of the coefficient of $y$ changes as $t_{1}$ changes its sign. Geometrically, this means that, for $t_{1}<0$, the real branch of $C^{(t)}$, which is tangent to ${\operatorname{Tor}}(\sigma)$ at $w$, lies in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$, while, for $t_{1}>0$, this real branch lies in $\overline{\mathcal{Q}}$, or vice versa (see Figure 1(c,d)).
At last, to have the family ${\mathcal{V}}$ smooth, representing the local deformation as above, and crossing the wall transversally, we need the following (global) transvrsality condition: the Zariski tangent space to the germ at $C^{*}$ of the family constrained by the given tangency conditions at fixed points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ and at two mobile points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ as well as by fixing the point $w_{0}$ has dimension zero. The routine computations, performed in preceding steps, reduce this condition to
$h^{0}\left({\mathbb{P}}^{1},{\mathcal{N}}_{{\Phi}^{\prime}}\left(-\sum_{p_{i}^{\sigma}\neq p}2k_{i}^{\sigma}p_{i}^{\sigma}+p_{i^{\prime}}^{\sigma^{\prime}}+p_{i^{\prime\prime}}^{\sigma^{\prime\prime}}-(2l_{1}+1)p^{\prime}-(2l_{2}+1)p^{\prime\prime}-p_{0}\right)\right)=0,$ (26)
where ${\Phi}^{\prime}:{\mathbb{P}}^{1}\to\widehat{C}\overset{{\Phi}}{\to}{\operatorname{Tor}}(P_{\Delta})$ is ${\Phi}$ combined with the normalization, $p^{\prime},p^{\prime\prime}$ are the preimages of $p\in\widehat{C}$. The vanishing follows from the fact that the degree of the line bundle is $-1$.
(6) Suppose that $\xi$ is as in Lemma 5.2(2vi) and that the path $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{|t-t^{*}|<\zeta}$ has the point $w_{0}(t)$ and two points $w_{1}(t),w_{2}(t)\in\widehat{\boldsymbol{w}}_{\partial}(t)$ in a fixed position, while the other two points $w_{3}(t),w_{4}(t)\in\widehat{\boldsymbol{w}}_{\partial}(t)$ are mobile. Also, for the dimension reason, we can assume that $w_{0}=w_{0}(t^{*})$ is in general position with respect to $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ on the rational curve $C$, which, in particular, yields the following: the point $p_{0}\in\widehat{C}$ is in general position with respect to the remaining points $p_{j}={\Phi}^{-1}(w_{j})\in\widehat{\boldsymbol{p}}_{\partial}\subset\widehat{C}$, $j=1,2,3,4$. By Lemma 2.4, the curve $C$ is immersed and smooth at $\widehat{\boldsymbol{w}}(t^{*})$. We can assume, in addition, that $C$ is nodal (the treatment of more complicated immersed singularities is very similar with a bit more involved notations).
Let $\xi(t)=[{\Phi}_{t}:(\widehat{C}_{t},{\boldsymbol{p}}_{t})\to{\operatorname{Tor}}(P_{\Delta})]\in\overline{\mathcal{M}}_{1,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}}(t))$, $|t-t^{*}|<\zeta$, run over the germ of a real one-dimensional subfamily of {¯$\{\overline{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))\}_{|t-t^{*}|<\zeta}$. It induces a one-dimensional family ${\mathcal{V}}\subset|{\mathcal{L}}_{P_{\Delta}}|$ of curves $C^{(t)}={\Phi}_{t,*}(\widehat{C}_{t})$ on ${\operatorname{Tor}}(P_{\Delta})$, where $C^{(t^{*})}={\Phi}_{*}(\widehat{C})=2C$. Any curve $C^{(t)}$, $t\neq t^{*}$, has four nodes in a neighborhood of each of the $(C^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta})[C])/2+1$ nodes of $C$. We show that the remaining $2c_{1}({\operatorname{Tor}}(P_{\Delta})[C]-4$ nodes of $C^{(t)}$ are located in a neighborhood of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$. The intersection of $C^{(t)}$ with a neighborhood of a point $w_{i}(t^{*})$, $1\leq i\leq 4$, is an immersed disc. Following the lines of part (2) of the proof of Lemma 4.3, introduce local conjugation-invariant coordinates $x,y$ in a neighborhood of $w_{j}(t^{*})$ such that $w_{i}(t)=(0,0)$ (which means that, if $w_{i}(t)$ is mobile, then the coordinate system moves too), the toric divisor containing $w_{i}(t)$ is ${\operatorname{Tor}}(\sigma)=\{y=0\}$, and ${\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})=\{y\geq 0\}$. Then, the equation $F_{t_{1}}(x)=0$ of $C_{t}$, where $t_{1}=t-t^{*}$, is given by formula (14), where we set $k=k_{i}$, and correspondingly, the geometry of $C^{(t)}$ in a neighborhood of $w_{i}(t^{*})$ is described by the real rational affine limit curve $C_{1}$ given by (15). Observe that $C_{1}$ has $k_{i}-1$ nodes and a global real one-dimensional branch having one vertical tangent. We then conclude that
  • the nodes of $C^{(t)}$ are located in a neighborhood of ${\operatorname{Sing}}\hskip 1.70709pt(C)\cup\widehat{\boldsymbol{w}}_{\partial}(t^{*})$;
  • the parameter $\gamma$ in (14) must be even, since otherwise, the change of sign of $t_{1}$ would lead to the change of the position of that global real branch with respect to the vertical tangent, but this position is determined by the double covering ${\Phi}:\widehat{C}\to C$; in particular, it follows that the change of sign of $t_{1}$ does not modify the contribution of the part of $C^{(t)}$ in a neighborhood of $w_{i}(t^{*})$ to formula (9) for $W_{1}(\xi_{t})$.
Let ${\operatorname{Span}}\{2C,C^{\prime}\}$ and ${\operatorname{Span}}\{C^{(t)},C^{\prime}_{t}\}$ be the lines in $|{\mathcal{L}}_{P_{\Delta}}|$ tangent to ${\mathcal{V}}$ at $2C$ and $C^{(t)}$, respectively. For $t\neq t^{*}$, the curve $C^{\prime}_{t}$ intersects $C^{(t)}$ at the point $w_{0}$, at each node, at each point $w_{j}^{\sigma}(t)$ with multiplicity $2k_{j}^{\sigma}$ or $12k_{j}^{\sigma}-$ 1 according as $w_{j}^{\sigma}$ is fixed or mobile, and at one more point $w^{\prime}$ (follows from the Bézout theorem). Lifting these points to $\widehat{C}^{(t)}$ and passing to the limit as $t\to t^{*}$, we obtain on $\widehat{C}$ the divisor
$D=p_{0}+4{\Phi}^{*}({\operatorname{Sing}}\hskip 1.70709pt(C))+(4k_{1}-2)p_{1}+(4k_{2}-2)p_{2}+(4k_{3}-3)p_{3}+(4k_{4}-3)p_{4}+p^{\prime}\ .$
Due to the general position of $w_{0}$ on $C$, we get $p^{\prime}\not\in\widehat{\boldsymbol{p}}_{\partial}\cup\{p_{0}\}$. We use this to show that for $t^{\prime}<t^{*}<t^{\prime\prime}$, the points $w_{i}(t^{\prime})$ and $w_{i}(t^{\prime\prime})$ are separated on the toric divisor by the point $w_{i}(t^{*})$, $i=3,4$. Indeed, otherwise, we would get $w_{i}(t^{\prime})=w_{i}(t^{\prime\prime})$ for, say, $i=3$ and for $t^{\prime}<t^{*}<t^{\prime\prime}$ depending on each other and converging to $t^{*}$. Lifting the intersection $C^{(t^{\prime})}\cap C^{(t^{\prime\prime})}$ upon $\widehat{C}_{t^{\prime}}$, we would obtain on $\widehat{C}$ the divisor
$D^{\prime}=p_{0}+4{\Phi}^{*}({\operatorname{Sing}}\hskip 1.70709pt(C))+(4k_{1}-2)p_{1}+(4k_{2}-2)p_{2}+(4k_{3}-2)p_{3}+(4k_{4}-3)p_{4}\ ,$
not linearly equivalent to $D$, which is a contradiction. Thus, the real part of the family ${\mathcal{V}}$ homeomorphically projects onto the interval $|t-t^{*}|<\zeta$, and the sign $W_{1}(\xi(t))$ remains constant as $0<|t-t^{*}|<\zeta$. $\Box$
6 Invariance in genus two: proof of Theorem 3.9
We closely follow the argument in the proof of Theorems 3.2 and 3.7. Take two sequences ${\boldsymbol{w}}(0)=({\boldsymbol{w}}_{\partial}(0),w_{1}(0),w_{2}(0))$ and ${\boldsymbol{w}}(1)=({\boldsymbol{w}}_{\partial}(1),w_{1}(1),w_{2}(t))$, satisfying the conditions of Theorem 3.7. We may assume that $\widehat{\boldsymbol{w}}_{\partial}(0),\widehat{\boldsymbol{w}}_{\partial}(1)\in\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ for some $\tau>0$. Then, we join these sequences by a generic path $\{\widehat{\boldsymbol{w}}_{\partial}(t)\}_{0\leq t\leq 1}$, in $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ and the points $w_{i}(0)$, $w_{i}(1)$ by a generic path $\{w_{i}(t)\}_{0\leq t\leq 1}$ in the respective open quadrant, $i=1,2$. Then we verify the constancy of $W_{2}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ inside the chambers (i.e., connected components of $\operatorname{Men}_{2}(\Delta)$) and in all possible wall-crossing events, for all $\kappa$.
6.1 Invariance inside chambers
Let us verify the constancy of $W_{2}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ along intervals $t^{\prime}<t<t^{\prime\prime}$ such that ${\boldsymbol{w}}(t)\in\operatorname{Men}_{2}(\Delta)$ for all $t^{\prime}<t<t^{\prime\prime}$. To this end, it is enough to show that the projection ${\mathcal{M}}_{2,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}\to(t^{\prime},t^{\prime\prime})$ has no critical points (cf. the proof of Theorems 3.2 and 3.7). This requirement is a consequence of the following statement similar to Lemmas 4.1 and 5.1 ( cf. [23, Lemma 2.3]).
Lemma 6.1
Let $\xi=[{\Phi}:(S_{2},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{2,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}})$, where $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{2}(\Delta)$. Then
$H^{0}\left(S_{2},{\mathcal{N}}_{\Phi}\left(-p_{1}-p_{2}-\sum_{s=0}^{2}\sum_{i=1}^{n_{s}}2k_{i}^{(s)}p_{i}^{(s)}+p_{n_{0}}^{(0)}\right)\right)=0,$ (27)
where $[{\Phi}:(S_{2},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in{\mathcal{M}}_{2,n}^{\mathbb{R}}({\mathbb{P}}^{2},\Delta,\widehat{\boldsymbol{w}}(t))$.
Proof. The non-vanishing in (27) would mean the existence of a real plane curve $C^{\prime}\neq C={\Phi}(S_{2})$ of degree $2d$ passing through $w_{1}$ and $w_{2}$, intersecting $D_{s}$, $s=0,1,2$, at each point $n_{i}^{(s)}$, $1\leq i\leq n_{s}$, $(s,i)\neq(0,n_{0})$, with multiplicity $2k_{i}^{(s)}$, intersecting $D_{0}$ at $w_{n_{0}}^{(0)}$ with multiplicity $12k_{n_{0}}^{(0)}-$ 1, and intersecting $C$ at each point $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)$ with multiplicity $\geq\delta(C,z)$. Note that $C^{\prime}$ must additionally intersect $C$ in an extra point on the real branch in $({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}$ and in an extra point on the real branch in ()±$({\mathbb{P}}_{\mathbb{R}})^{2}_{\pm}$. However, this leads to a contradiction with the Bézout theorem:
$4d^{2}\geq 4+\sum_{s=0}^{2}\sum_{i=1}^{n_{s}}2k_{i}^{(s)}-1+\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}2\delta(C,z)$
$=4+(6d-1)+((2d-1)(2d-2)-4)=4d^{2}+1.$
$\Box$
6.2 Walls in the space of constraints
Let $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in\overline{\mathcal{M}}_{2,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}}(t))_{0\leq t\leq 1}$. In this section, we assume that $\widehat{C}$ does not contain contractible components, i.e., we just contract them (if any) and we call the singularities of $\widehat{C}$ “nodes”. Accordingly, we allow the points of $\widehat{\boldsymbol{p}}$ to collate. Note that such a contraction does not affect the arithmetic genus, since a generic curve $\widehat{C}=S_{2}$ has three real ovals.
Lemma 6.2
(1) The following elements $\xi=[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]$ cannot occur in ¯$\overline{\mathcal{M}}_{2,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}}(t))_{0\leq t\leq 1}$:
  1. (1i)
    $\widehat{C}$ is a reducible connected curve of arithmetic genus $2$ with a component mapped onto a toric divisor;
  2. (1ii)
    $\widehat{C}$ is a connected curve of arithmetic genus $2$ with at least two nodes;
  3. (1iii)
    a component of $\widehat{C}$ multiply covers its image.
(2) If $\xi=[{\Phi}:(\widehat{C},\widehat{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in\overline{\mathcal{M}}_{2,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ when $\widehat{\boldsymbol{w}}(t^{*})$ does not satisfy the conclusions of Lemma 2.4, then $\xi$ is of one of the following types:
  1. (2i)
    $\widehat{C}$ is a smooth curve of genus $2$, $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points, ${\Phi}$ is birational onto its image, but not an immersion; furthermore, the germs ${\Phi}(\widehat{C},p_{i})$, $i=1,2$, and the germs ${\Phi}(\widehat{C},p_{i}^{(s)})$ for all $s=0,1,2$ and $k_{i}^{(s)}\geq 2$ are smooth;
  2. (2ii)
    $\widehat{C}$ is a smooth curve of genus $2$, $\widehat{\boldsymbol{w}}_{\partial}$ contains $n-2$ distinct points so that $w_{i}^{(s)}(t^{*})=w_{j}^{(s)}(t^{*})$ for some $s=0,1,2$ and $i\neq j$, and $p_{i}^{(s)}=p_{j}^{(s)}=p\in{\boldsymbol{p}}$; in addition, the map ${\Phi}$ is an immersion;
  3. (2iii)
    $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $g(\widehat{C}_{1})+g(\widehat{C}_{2})=2$, and the only node joining $\widehat{C}_{1}$ and $\widehat{C}_{2}$ is mapped to (×$({\mathbb{R}}^{\times})^{2}\subset{\mathbb{P}}_{\mathbb{R}}^{2}$; $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points; the maps ${\Phi}:\widehat{C}_{i}\to{\mathbb{P}}^{2}$, $i=1,2$, are birational immersions;
  4. (2iv)
    $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $g(\widehat{C}_{1})+g(\widehat{C}_{2})=2$; $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points; the only node $p$ joining $\widehat{C}_{1}$ and $\widehat{C}_{2}$ is mapped to some point $w_{i}^{(s)}(t^{*})\in\widehat{\boldsymbol{w}}(t^{*})$; the maps ${\Phi}:\widehat{C}_{i}\to{\mathbb{P}}^{2}$, $i=1,2$, are birational immersions; the germ ${\Phi}(\widehat{C}_{j},p)$ intersects $D_{s}$ with multiplicity $2l_{j}$, $j=1,2$, where $l_{1}+l_{2}=k_{i}^{(s)}$ and these two germs intersect each other with multiplicity $\min\{2l_{1},2l_{2}\}$;
  5. (2v)
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points, $\widehat{C}$ is an irreducible elliptic curve with one node $p$, the map ${\Phi}$ is an immersion that sends $p$ to some point $w_{i}^{(s)}(t^{*})$, center of two smooth branches intersecting $D_{s}$ with odd multiplicities $12l_{1}+1,2l_{2}+$ 1, where $l_{1}+l_{2}+1=k_{i}^{(s)}$, and intersecting each other with multiplicity $\min\{2l_{1}+1,2l_{2}+1\}$; furthermore, each of the other points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ is a center of one smooth branch.
Proof. (1i) We can assume that the sequences $\{{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ avoid $\tau_{0}$-neighborhoods of the intersection points of the coordinate axes, for some $\tau_{0}>0$. Assume that ${\Phi}(\widehat{C})$ contains some coordinate axis.
Observe that if ${\Phi}(\widehat{C})$ contains at least one coordinate axis, then it contain all of them. Indeed, otherwise we would have ${\boldsymbol{w}}(t)$ arbitrarily close to the intersection points of the coordinate axes, contrary to the above assumption. Furthermore, the intersection points of the coordinate axes must lift to nodes of $\widehat{C}$, and they smooth out along the path $\{{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$. In particular, the union $\widehat{C}_{1}$ of the components of $\widehat{C}$ mapped onto the coordinate axes is a curve of arithmetic genus $\geq 1$.
The union $\widehat{C}_{2}$ of the other components of $\widehat{C}$ cannot be mapped onto a plane curve of odd degree. Indeed, otherwise, one would get an odd number of real branches of ${\Phi}(\widehat{C}_{2})$ intersecting $D_{0}$ with odd multiplicity yielding a one-dimensional real arc in the quadrant $\{x_{1}/x_{0}<0,\ x_{2}/x_{0}<0\}$ which is not possible. Thus, ${\Phi}_{*}(\widehat{C}_{2})$ has an even degree. If $\widehat{C}_{2}$ has an arithmetic genus 00, then the real part of ${\Phi}(\widehat{C}_{2})$ contains a connected graph passing through $w_{1}$ and $w_{2}$, and then there should be four real local branches of ${\Phi}(\widehat{C}_{2})$ centered on $D_{1,+}\cup D_{2,+}$ that intersect $D_{1}\cup D_{2}$ with odd multiplicity. Hence, the centers of these branches must lift to the nodes of $\widehat{C}$, which would yield that the arithmetic genus of $\widehat{C}$ is at least $4$ a contradiction. Thus, $\widehat{C}_{1}$ and $\widehat{C}_{2}$ both have arithmetic genus $1$, and are joined by one node. It is easy to see, that this node must be mapped by ${\Phi}$ to the center of the unique local branch of ${\Phi}(\widehat{C}_{2})$ on $D_{0,+}$. This local branch must lie in the closure of $({\mathbb{P}}_{\mathbb{R}}^{2})^{+}$. Hence the image of one connected component of the real part of $\widehat{C}_{2}$ must pass through, say, $w_{1}$ and contain the real local branch centered on $D_{0,+}$. It follows that ${\Phi}(\widehat{C}_{2})$ has two real local branches centered on $D_{1,=}$ and intersecting $D_{1}$ with odd multiplicity. Thus, their centers must lift to two nodes of $\widehat{C}$ joining $\widehat{C}_{1}$ and $\widehat{C}_{2}$ making the arithmetic genus of $\widehat{C}$ at least $3$ which is a contradiction.
(1ii) We argue on contradiction assuming that $\widehat{C}$ has $k\geq 2$ nodes and no component of $\widehat{C}$ is mapped onto a coordinate axis.
First, $\widehat{C}$ cannot have $\geq 3$ components, since it would impose two extra Menelaus conditions on the disposition of ${\boldsymbol{w}}_{\partial}$ defining a stratum of codimension $2$ in ${\mathcal{W}}$ which must be avoided by the path $\{{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$.
Second, $\widehat{C}$ cannot be an irreducible rational curve. Indeed, if ${\boldsymbol{w}}_{\partial}\in{\mathcal{W}}_{\partial}$ is a general position, then ${\boldsymbol{w}}(\widehat{C})$ belongs to a finite set, see Lemma 2.4. Hence, if ${\boldsymbol{w}}_{\partial}\in{\mathcal{W}}_{\partial}$ varies in a stratum of codimension $1$, then we obtain a one-dimensional family of rational curves, and this imposes an extra condition on the position of the points $w_{1},w_{2}$, altogether defining a stratum of codimension $\geq 2$ in ${\mathcal{W}}$.
Third, $\widehat{C}$ cannot be the union of a rational and an elliptic component. Then we obtain an extra Menelaus condition on ${\boldsymbol{w}}_{\partial}$, while the parts of ${\boldsymbol{w}}_{\partial}$ lying on the image of the rational and of the elliptic components are in general position. Thus, ${\boldsymbol{w}}_{\partial}$ and one of the points $w_{1},w_{2}$ determine a finite set of pairs of a rational and an elliptic components of $\widehat{C}$ (Lemma 2.4), which imposes a condition on the location of the second point in $\{w_{1},w_{2}\}$, hence, defining a stratum of codimension $\geq 2$ in ${\mathcal{W}}$.
(1iii) Suppose that $\widehat{C}$ is irreducible, ${\Phi}:\widehat{C}\twoheadrightarrow C\hookrightarrow{\mathbb{P}}^{2}$ a multiple covering. By the Riemann-Hurwitz formula, $C$ cannot be of geometric genus $2$. Also, $C$ cannot be rational since otherwise, the fixed position of ${\boldsymbol{w}}_{\partial}$ determines a finite set of rational curves, and hence two conditions are imposed on the position of $w_{1},w_{2}$. This defines a stratum of codimension $\geq 2$ in ${\mathcal{W}}$ avoided by the path $\{{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$. The only remaining option is an elliptic curve $C$. As in the case of the rational $C$, we derive that ${\boldsymbol{w}}_{\partial}$ varies in an open dense subset of ${\mathcal{W}}_{\partial}$, while the position of $w_{1},w_{2}$ is constrained by one condition. It follows that all the points of the sequence $\widehat{\boldsymbol{w}}_{\partial}$ are distinct, and hence they all must be ramification points of the covering. However, #$\#\widehat{\boldsymbol{w}}_{\partial}\geq 3$ which yields a contradiction in the Riemann-Hurwitz formula.
In view of (1ii), it remains to study the case of two-component $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$ such that either $g(\widehat{C}_{1})=2$, $g(\widehat{C}_{2}=0$, or $g(\widehat{C}_{1})=\widehat{C}_{2})=1$. It follows that ${\boldsymbol{w}}_{\partial}$ is constrained by an extra Menelaus condition, and this is the only condition imposed. Hence, all the points of $\widehat{\boldsymbol{w}}_{\partial}$ differ from each other, and the genus of the image of any component must be equal to the genus of that component. The argument of the preceding paragraph leaves the only option of a multiple covering ${\Phi}:\widehat{C}_{2}\twoheadrightarrow C_{2}\hookrightarrow{\mathbb{P}}^{2}$ with rational $\widehat{C}_{2}$ and $C_{2}$. If $m$ is the degree of the covering, then each point $w_{i}^{(s)}\in C_{2}$ is a ramification point of the covering of index $m-1$, and hence we get a contradiction in the Riemann-Hurwitz formula:
$2m-(m-1)\cdot\#(\widehat{\boldsymbol{w}}_{\partial}\cap C_{2})\geq 2,\quad\#(\widehat{\boldsymbol{w}}_{\partial}\cap C_{2})\geq 3$
(2) The classification of elements of
¯$\overline{\mathcal{M}}_{2,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))\setminus{\mathcal{M}}_{2,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$
is derived literally in the same manner as Lemma 5.2(2ii,2iii,2iv,2v0). We only notice in addition that the cases including multiple coverings do not occur as we showed in part (1iii). $\Box$
6.3 Wall-crossing
In the analysis of wall-crossings, we use arguments similar to the ones presented in Sections 4 and 5. The following lemma completes the proof of Theorem 3.9.
Lemma 6.3
The number $W^{\kappa}_{2}(\Delta,\widehat{\boldsymbol{w}}(t))$ remains constant in each of the wall-crossing events described in Lemma 6.2(2).
Proof. All local considerations in the proof of Theorems 3.2 and 3.7 can literally be transferred to our situation. Thus, we only confirm the global transversality relations which all take form of certain cohomology vanishing conditions.
(1) The wall-crossing described in Lemma 6.2(2i) coincides with that in Lemma 5.2(2ii). The transversality condition in the treatment of this wall-crossing reads (cf. formula (23))
$H^{0}(S_{2},{\mathcal{O}}_{S_{2}}({\boldsymbol{d}}-p_{1}-p_{2}))=0,\quad{\boldsymbol{d}}\in{\operatorname{Div}}(S_{2}),$ (28)
where in the notation of the proof of Lemma 5.3, part (2),
$\deg{\boldsymbol{d}}=C^{2}-\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}2\delta(C,z)-\sum_{w\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})}\dim{\mathcal{O}}_{C,w}/I_{w}$
$=4d^{2}-(4d^{2}-6d-2)-(6d-1)=3>(2g-2)\big{|}_{g=2}=2.$
Thus, the Riemann-Roch formula yields $h^{0}(S_{2},{\mathcal{O}}_{S_{2}}({\boldsymbol{d}}))=2$, and hence (28) holds due to the general position of $w_{1},w_{2}$ on $C={\Phi}(S_{2})$.
(2) The wall-crossing described in Lemma 6.2(2ii) coincides with that in Lemma 5.2(2iii) (the case of $C$ unibranch at $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})$, i.e., $p_{i}^{\sigma}=p_{j}^{\sigma}$). The required transversality condition coincides with relation (28) confirmed in the preceding paragraph.
(3) The wall-crossing described in Lemma 6.2(2iii) coincides with that in Lemma 5.2(2iv), the subcase of ${\Phi}:\widehat{C}_{i}\to{\operatorname{Tor}}(\Delta)$, $i=1,2$, birational immersions. The local study is a word-for-word copy of the proof of Lemma 4.3, part (5a) (restricted to the case of birational immersions). Similarly to the proof of Lemma 4.3, part (5b), and the proof of Lemma 5.3, part (4), the global transversality condition amounts to
$h^{0}(\widehat{C}_{1},{\mathcal{O}}_{\widehat{C}_{1}}({\boldsymbol{d}}_{1}-\{p_{1},p_{2}\}\cap\widehat{C}_{1}))=h^{0}(\widehat{C}_{2},{\mathcal{O}}_{\widehat{C}_{2}}({\boldsymbol{d}}_{1}-\{p_{1},p_{2}\}\cap\widehat{C}_{2}))=0,$
where
$\deg{\boldsymbol{d}}_{1}=2g(\widehat{C}_{1})-1,\quad\deg{\boldsymbol{d}}_{2}=2g(\widehat{C}_{2})-1.$
To verify the above $h^{0}$-vanishing conditions we argue as in step (2) above.
(4) The last wall-crossing to consider is associated with the wall described in Lemma 6.2(2v). That wall is very similar to the wall appearing in Lemma 5.2(2v)]. The local study of the wall-crossing is a word-for-word copy of the study performed in the proof of Lemma 5.3, part (5). Thus, we only address the global transversality condition analogous to that in formula (26):
$h^{0}({\mathbb{E}},{\mathcal{N}}_{{\Phi}^{\prime}}({\boldsymbol{d}}))=0,$
${\boldsymbol{d}}=-\sum_{p_{i}^{(s)}\neq p}2k_{i}^{(s)}p_{i}^{(s)}+p_{i^{\prime}}^{(s^{\prime})}+p_{i^{\prime\prime}}^{(s^{\prime\prime})}-(2l_{1}+1)p^{\prime}-(2l_{2}+1)p^{\prime\prime}-p_{1}-p_{2},$
where ${\Phi}^{\prime}:{\mathbb{E}}\to\widehat{C}\overset{{\Phi}}{\to}{\mathbb{P}}^{2}$ is ${\Phi}$ combined with the normalization, ${\mathcal{N}}_{{\Phi}^{\prime}}$ is the normal bundle, $p_{i^{\prime}}^{(s^{\prime})},p_{i^{\prime\prime}}^{(s^{\prime\prime})}$ are mobile points in $\widehat{\boldsymbol{w}}_{\partial}$ (while the other points are fixed), and $p^{\prime},p^{\prime\prime}$ are the preimages of the node $p$ in ${\mathbb{E}}$. It is easy to see that
$\deg{\boldsymbol{d}}=0.$
However, the points $w_{1},w_{2}$ are in general position on $C={\Phi}(\widehat{C})$, and hence $p_{1},p_{2}$ are in general position on ${\mathbb{E}}$, which by the Riemann-Roch formula implies the desired $h^{0}$-vanishing. $\Box$